### v1

# Robust $Q$-Learning for finite ambiguity sets

Abstract.
In this paper we propose a novel $Q$-learning algorithm allowing to solve distributionally robust Markov decision problems for which the ambiguity set of probability measures can be chosen arbitrarily as long as it comprises only a finite amount of measures. Therefore, our approach goes beyond the well-studied cases involving ambiguity sets of balls around some reference measure with the distance to reference measure being measured with respect to the Wasserstein distance or the Kullback–Leibler divergence. Hence, our approach allows the applicant to create ambiguity sets better tailored to her needs and to solve the associated robust Markov decision problem via a $Q$-learning algorithm whose convergence is guaranteed by our main result. Moreover, we showcase in several numerical experiments the tractability of our approach.
Keywords: Q-Learning, Markov Decision Processes, Model Uncertainty
July 5, 2024
1National University of Singapore, Department of Mathematics,
21 Lower Kent Ridge Road, 119077.
Corresponding author, email:jul__\__ses@nus.edu.sg
1. Introduction
Markov decision problems are discrete time stochastic control problems involving an agent who can influence the environment by executing actions affecting the distribution of the realization of the subsequent state of the environment. In turn, depending on the state realization, the agent receives a feedback in form of a reward and decides on her next action. The objective of the agent is then to optimize the cumulative expected rewards obtained over an infinite time horizon by choosing the highest rewarding sequence of actions.
By construction, the described optimization problem depends crucially on the involved distributions for the state transitions. In practice, frequently in model-based approaches these distributions are estimated ([1], [32], [37]) from data or implicitly learned on-line in model-free approaches (see, e.g., [8], [39]). However chosen, the underlying model is highly prone to model misspecification, i.e., if the underlying distribution changes after training, the agent is trained to the wrong underlying distribution and successful execution of actions (i.e., receiving high rewards) typically will fail if the distribution change is too large. A solution to this problem is to take distributional uncertainty into account, i.e., to train the agent already by using an ambiguity set of probability distributions and train the agent according to the worst case measure contained in the ambiguity set, following the paradigm that during execution the distribution of the state transition might change but still is contained in a pre-defined ambiguity set. In this paper we follow this paradigm and propose a novel numerical algorithm to solve Markov decision problems while accounting for distributional uncertainty by considering arbitrary finite ambiguity sets.
1.1. Our Contribution
One frequently used numerical method to solve (non-robust) Markov decision problems that has turned out to work efficiently in practice is the so-called $Q$-learning algorithm which initially was introduced by Watkins in his PhD thesis, see [46], as well as [47] for the accompanying paper. For a general introduction to $Q$-learning algorithms we refer to [11] and [19], and for more example of associated applications, see among many others [3], [4], [9], [10], [16], [21], and [26].
Our contribution is to present a novel $Q$-learning algorithm that allows to take distributional uncertainty for the state transitions of Markov decision problems into account and which therefore respects the so called Knightian uncertainty ([20]), i.e., the uncertainty of having chosen the correct model.
We respect this model uncertainty by, instead of fixing a unique probability distribution for the state transitions, allowing the state transition between two subsequent states to be realized according to any distribution from a set of predetermined transition probabilities and then to consider the worst-case measure to compute rewards. Moreover, to account for the fact that transition probabilities may change over time, we consider a time-inhomogeneous formulation, i.e., we allow different transition probabilities in every instance of time as long as the transition probabilities are contained in the ambiguity set. To solve this robust formulation of a Markov decision process (as introduced in [28]) under the worst-case measure we develop a novel $Q$-learning algorithm.
The finite ambiguity set of transition probabilities can be designed by the applicant and its shape is, in particular, not restricted to measures that are in a certain sense close to a reference measure. This allows to consider different ambiguity sets as it were possible in the case of ambiguity sets that are defined as balls around reference measures. Put differently, using our approach, an applicant can simply include any possible scenarios in which she is interested and hence is able to control the shape and content of the ambiguity sets and hence the dynamics of the underlying Markov decision process that are deemed to be admissible. Moreover, in Section 5 we provide a proof guaranteeing convergence of our algorithm.
1.2. Related Work
While the study of distributionally robust optimization (DRO) problems and distributionally robust Markov decision problems and problems has been an active research topic already in the past decade (compare, e.g., [5], [6], [15], [23], [28], [29], [30], [34], [35], [41], [45], [48], [49], [50], and [51]), the discussion and construction of $Q$-learning algorithms to solve these sequential decision making problems numerically has only become an active research topic very recently.
In [6], [22], and [44] Q-learning algorithms for ambiguity sets that are constructed as balls around some reference measures with respect to the Kullback–Leibler divergence are discussed.
Moreover, in [27], the authors present a novel $Q$-learning algorithm for the specific case that the ambiguity set is given by a Wasserstein-ball around a reference measure.
To the best of our knowledge going beyond the aforementioned cases imposes a novelty, and has not been studied yet.
1.3. Structure
The remainder of this paper is as follows. In Section 2 we present the setting and the associated stochastic control problem. A numerical solution to this optimization problem via $Q$-learning is presented in Section 3 together with our main result guaranteeing convergence of the $Q$-learning algorithm. In Section 4 we provide numerical experiments showcasing the tractability of our $Q$-learning algorithm. The proofs are reported in Section 5.
2. Setting and specification of the problem
In this section we introduce the underlying framework which we use to establish a robust Q-learning algorithm for finite ambiguity sets.
2.1. Setting
Optimal control problems, such as Markov decision problems, are formulated using a state space comprising all possible states accessible by an underlying stochastic process. We model this state space by a $d$-dimensional finite Euclidean subset $\mathcal{X}\subset\mathbb{R}^{d}$ and eventually aim at solving a robust control problem over an infinite time horizon. Hence, we define the state space over the entire time horizon via
$\Omega:=\mathcal{X}^{\mathbb{N}}=\mathcal{X}\times\mathcal{X}\times...$
equipped with the corresponding $\sigma-$algebra $\mathcal{F}:=2^{\mathcal{X}}\otimes 2^{\mathcal{X}}\otimes\cdots$. Next, let $(X_{t})_{t\in\mathbb{N}}$ be the state process, i.e., the stochastic process on $\Omega$ describing the states attained over time. We denote the finite set of possible actions by $A\subset\mathbb{R}^{m}$, where $m\in\mathbb{N}$ represents the dimension of the action space. The set of admissible policies is then given by
$\mathcal{A}:=$$\{\textbf{a}=(a_{t})_{t\in\mathbb{N}}\ |\ (a_{t})_{t\in\mathbb{N}}:\Omega\rightarrow A;\ a_{t}\text{ is }\ \sigma(X_{t})-\text{measurable}\ \text{for all }t\in\mathbb{N}\}$
$=$ $\{(a_{t}(X_{t}))_{t\in\mathbb{N}}\ |\ a_{t}:\mathcal{X}\rightarrow A\ \text{Borel\ measurable }\text{for all }t\in\mathbb{N}\}.$
[…]
(2.1) $\mathcal{X}\times A$ $\rightarrow(\mathcal{M}_{1}(\mathcal{X})^{N},\tau_{0})$
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\},\ \ \ \ \ \ N\in\mathbb{N},$
where for all $k\in\{1,\dots,N\}$ and for all $(x,a)\in\mathcal{X}\times A$ we have that $\mathbb{P}^{(k)}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. Note that in the degenerate case where $\mathcal{P}(x,a)$ contains only a single probability distribution, we get a classical Markov decision process (i.e. a non-robust one), compare, e.g., [7].
The ambiguity set of admissible probability distributions on $\Omega$ depends on the initial state $x\in\mathcal{X}$ and the policy $\textbf{a}\in\mathcal{A}$. Let $\delta_{x}\in\mathcal{M}_{1}(\mathcal{X})$ denote the Dirac measure at point $x\in\mathcal{X}$. Then, we define for every state-action pair $(x,\textbf{a})\in\mathcal{X}\times\mathcal{A}$ the underlying set of admissible probability measures of the stochastic process $(X_{t})_{t\in\mathbb{N}}$ by
[…]
$\text{and }\mathbb{P}_{t}\in\mathcal{P}(x_{t},a_{t}(x_{t})),\ \text{ for all }x_{t}\in\mathcal{X}\Bigl{\}},$
where the notation $\mathbb{P}=\delta_{x}\otimes\mathbb{P}_{0}\otimes\mathbb{P}_{1}\otimes...\in\mathfrak{P}_{x,a}$ abbreviates
$\mathbb{P}(B)=\sum_{x_{0}\in\mathcal{X}}...\sum_{x_{t}\in\mathcal{X}}...1\mathrm{l}_{B}((x_{t})_{t\in\mathbb{N}})...\mathbb{P}_{t-1}(x_{t-1};\{x_{t}\})...\mathbb{P}_{0}(x_{0};\{x_{1}\})\delta_{x}(\{x_{0}\}),\ \ \ \ \ \ B\in\mathcal{F}.$
The construction of ambiguity sets $\mathfrak{P}_{x,\textbf{a}}$ provided in (2.1) and (2.2) is in the related literature often referred to as a $(s,a)$-rectangular ambiguity set, (compare [17] and [48]) and enables to obtain a dynamic programming principle of the associated Markov decision problem.
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$.
2.2. Finite ambiguity sets and other ambiguity sets
Note that the paradigm under which we work in this paper is fundamentally different from the approaches pursued, e.g., in [6], [22], and [27] where the corresponding ambiguity sets of probability measures are defined as balls around some reference measure.
The Q-learning approaches presented in [6], [22], and [27] provide solutions to account for a potential misspecification of a reference measure, by allowing deviations from it where the size of the deviations is measured by the respective distances (Kullback–Leibler distance and Wasserstein distance). However, these approaches allow not to control of which type the distributions in the ambiguity set are but simply consider all measures that are in a certain sense close to the reference measure. If an applicant is instead interested in controlling the distributions contained in the ambiguity set, for example, by allowing only for a specific type of parametric distributions with a finite set of possible parameters or by considering an asymmetric ambiguity set, the applicant can use the $Q$-learning approach presented in this paper. Another relevant situation is the consideration of multiple estimated models / probability measures obtained for exampled via different estimation methods. To account for all of these estimations, an applicant can simply construct an ambiguity set containing all estimated probability measures.
Beyond these use cases, our approach also allows to combine different types of distributions, if desired.
The main difference therefore is that an applicant of the approach presented in this paper can control exactly what types of distributions are deemed possible whereas the approaches from [6], [22], and [27] account for a misspecification of one single reference measure.
2.3. Specification of the optimization problem
The robust optimization problem consists, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in\mathbb{R}$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
(2.3) $\mathcal{X}\ni x\mapsto V(x):=\sup_{\mathbf{a}\in\mathcal{A}}\inf_{\mathbb{P}\in\mathfrak{P}_{x,\textbf{a}}}\mathbb{E}_{\mathbb{P}}\left[\sum^{\infty}_{t=0}\alpha^{t}\cdot r(X_{t},a_{t}(X_{t}),X_{t+1})\right]$
describes the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ and when executing the optimal policy $\textbf{a}\in\mathcal{A}$ after having started in initial state $x\in\mathcal{X}$.
2.4. Dynamic Programming
We consider the one step optimization problem
(2.4) $\mathcal{T}V(x):=\max_{a\in A}\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],\ \ \ \ x\in\mathcal{X}$
where $\mathcal{X}\ni x\mapsto V(x)$ is the value function defined in (2.3). Analogue to the definition of the non-robust $Q$-value function, see e.g. [46], we define the robust optimal Q-value function by
(2.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q^{*}(x,a):=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
We can interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $V(x)$. By using the above definitions we indeed obtain the following dynamic programming equation, also famously known in the non-robust formulation as Bellman equation, compare e.g. [12, p.164] for a non-robust version.
Proposition 2.1.
Assume $10<\alpha<$ 1. Then, for all $x\in\mathcal{X}$ we have
$\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)=V(x).$
3. Robust Q-Learning Algorithm
In this section we propose with Algorithm 1 a novel distributionally robust $Q$-learning algorithm and provide in Theorem 3.1 conditions for its convergence.
Algorithm 1 Robust $Q$-learning for finite ambiguity sets
  Input:
   State space $\mathcal{X}\subseteq\mathbb{R}^{d}$;  Initial state $x_{0}\in\mathcal{X}$;  Action space $A\subseteq\mathbb{R}^{m}$;
   Reward function $r$;  Discount factor $\alpha\in(0,1)$;
   Ambiguity set $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\ldots,\mathbb{P}^{(N)}(x,a)\right\}$;
   Policy $(a_{t})_{t}\in\mathcal{A}$;  Sequence of learning rates $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$;
  Initialize $Q_{0}(x,a)$ for all $(x,a)\in\mathcal{X}\times A$ to an arbitrary real value;
  Set $X_{0}=x_{0}$;
  for $t=0,1,\ldots$ do
     Determine an index $k_{t}^{*}\in\{1,\ldots,N\}$ by
(3.1) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right];$
     Sample $X_{t+1}\sim\mathbb{P}^{(k_{t}^{*})}\left(X_{t},a_{t}(X_{t})\right)$;
     for all $(x,a)\in\mathcal{X}\times A$ do
        Update $Q_{t}$ via
(3.2) $Q_{t+1}(x,a)\leftarrow\begin{cases}(1-\widetilde{\gamma}_{t})\cdot Q_{t}(x,a)+\widetilde{\gamma}_{t}\cdot\left(r(x,a,X_{t+1})+\alpha\max\limits_{b\in A}Q_{t}(X_{t+1},b)\right)&\text{if }(x,a)=\left(X_{t},a_{t}(X_{t})\right),\\ Q_{t}(x,a);&\text{else}\end{cases}$
     end for
  end for
  Output: A sequence $(Q_{t}(x,a))_{t\in\mathbb{N},x\in\mathcal{X},a\in A}$;
Given a policy $\textbf{a}\in\mathcal{A}$ and some initial state $x_{0}\in\mathcal{X}$ we define a proability measure by
(3.3) $\mathbb{P}_{x_{0},\textbf{a}}:=\delta_{x_{0}}\otimes\mathbb{P}_{0}^{(k_{0}^{*})}(\cdot,a_{0}(\cdot))\otimes\mathbb{P}_{1}^{(k_{1}^{*})}(\cdot,a_{1}(\cdot))\otimes\cdots\in\mathfrak{P}_{x_{0},\textbf{a}},$
where $k_{t}^{*}$ is as defined in (3.1). Note that the probability measure defined in  (3.3) is the probability measure according to which the states are sampled when applying Algorithm 1.
Our following main result now shows that the function $(Q_{t})_{t\in\mathbb{N}}$ obtained as the output of Algorithm 1 converges pointwise almost surely towards the optimal robust $Q$-value function $Q^{*}$ defined in (2.5).
Theorem 3.1.
Let $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$, and define for all $(x,a)\in\mathcal{X}\times A$ and for all $t\in\mathbb{N}$
(3.4) $\gamma_{t}(x,a,X_{t}):=1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{t}.$
Moreover, let $10<\alpha<$ 1 and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(3.5) $\sum^{\infty}_{t=0}\gamma_{t}(x,a,X_{t})=\infty,\ \sum^{\infty}_{t=0}\gamma_{t}^{2}(x,a,X_{t})<\infty,\ \text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
[…]
(3.6) $\lim_{t\rightarrow\infty}Q_{t}(x,a)=Q^{*}(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Remark 3.2 (On the implementation).
  • (a)
    Note that in many situations one can improve the convergence speed by replacing $\widetilde{\gamma}_{t}$ with $\widetilde{\gamma}_{\operatorname{visits}_{t}(x,a)}$, where we initially set $\operatorname{visits}_{0}\equiv 0$, and then for all $(x,a)\in\mathcal{X}\times A$ and $t=1,2,\dots,$
    $\operatorname{visits}_{t}(x,a)\leftarrow\left\{\begin{array}[]{ll}\operatorname{visits}_{t}(x,a)+1&\text{if}\ (x,a)=\left(X_{t},~{}a_{t}(X_{t})\right),\\ \operatorname{visits}_{t}(x,a)&\text{else},\end{array}\right.$
    compare, e.g., [24]. The idea behind this modification is to account for the exploration-exploitation trade off by reducing the learning rate the more often we have visited a state and hence the more confident the algorithm is w.r.t.  choosing the correct optimal action, and conversely to apply a larger learning rate if the state-action pair was visited less often until time $t$.
  • (b)
    Note that applying Algorithm 1 requires to choose a pre-determined policy $\textbf{a}\in\mathcal{A}$. Implied from usual practice in non-robust $Q$-learning (([13], Chapter 9), [25], or [40]) and robust $Q$-learning [27], a reasonable choice seems to be the $\varepsilon$-greedy policy defined by
    $\mathcal{X}\ni x\mapsto a_{t}(x):=\begin{cases}\operatorname{argmax}_{b\in B}Q_{t}(x,b)&\text{with probability }1-\epsilon_{\operatorname{greedy}},\\ a\sim\mathcal{U}(A)&\text{with probability }\epsilon_{\operatorname{greedy}},\end{cases}$
    for some $\epsilon_{\operatorname{greedy}}>0$, and where $a\sim\mathcal{U}(A)$ means that a random action $a$ is chosen uniformly at random from the finite set $A$.
  • (c)
    Due to Equation (3.1), increasing the number of probability measures $N$ linearly increases also the computation time linearly as bloating the ambiguity set by a factor $N$ corresponds to computing $N$ times as many expectations of the form $\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)]$ in order to determine $k_{t}^{*}$.
4. Examples
In this section we provide two numerical examples comparing Algorithm 1 with other robust and non-robust $Q$-learning algorithms. To apply the numerical method from Algorithm 1, we use for all of the following examples a sequence of learning rates defined by $\widetilde{\gamma}_{t}=\frac{1}{1+t}$ for $t\in\mathbb{N}$, a discount factor of $\alpha=0.95$, as well as an $\epsilon$-greedy policy with $\epsilon_{\operatorname{greedy}}=0.1$. Finally, we run [^2] We consider the algorithm as being converged if further iterations do not change the optimal action derived from $Q$-learning, i.e. if $\operatorname*{arg\,max}_{b\in A}Q_{t}(x,b)$ remains constant for further iterations $t$ for all $x\in\mathcal{X}$. This was the case for all presented experiments. Algorithm 1 with $1~{}000~{}000$ iterations on a processor: 13th Gen Intel(R) Core(TM) i7-13700KF 3.40 GHz. Further details of the implementation can be found under https://github.com/CecileDecker/FiniteQLearning.git.
4.1. Coin Toss
As in [28, Example 4.1] we consider an agent playing a coin toss game.
At each time step $t\in\mathbb{N}$ the agent observes the result of $10$ coins where an outcome of heads corresponds to $1$, and tails corresponds 00. The state $X_{t}$ at time $t\in\mathbb{N}$ is then given by the sum of the $10$ coins value, i.e., we have $\mathcal{X}:=\{0,...,10\}$.
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e. $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e. $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\$$, however if the agent is wrong she has to pay $1\$$. The agent also has the possibility not to play. We model this by considering the following reward function:
(4.1) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a1\mathrm{l}_{\{x<x^{\prime}\}}-a1\mathrm{l}_{\{x>x^{\prime}\}}-|a|1\mathrm{l}_{\{x=x^{\prime}\}},$
where the possible actions are given by $A:=\{-1,0,1\}$ with $a=1$ corresponding to betting $X_{t+1}>X_{t}$, $a=0$ to not playing, and $a=-1$ to betting $X_{t+1}<X_{t}$. In a next step, we define two different ambiguity sets via [^3] We denote by $\operatorname*{Bin}(n,p)$ a binomial distribution with $n$ number of trials and $p$ being the probability of success in a single trial.
(4.2) $\mathcal{P}^{1}(x,a):=$$\{\operatorname*{Bin}(10,0.5),~{}\operatorname*{Bin}(10,0.6)\},$
(4.3) $\mathcal{P}^{2}(x,a):=$$\{\operatorname*{Bin}(10,0.5),~{}\operatorname*{Bin}(10,0.3)\},$
and we aim at comparing Algorithm 1 with the Algorithm from [28] which however does not allow to build the same assymmetric sets as in (4.2) and (4.3). To build comparative ambiguity sets with Wasserstein uncertainty we define ambiguity sets such that all probability measures from $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ are contained, respectively. This leads to [^4] Here $W_{1}$ denotes the Wasserstein-$1$ distance, compare for more details, e.g., [43].
(4.4) $\mathcal{P}^{3}(x,a):=$$\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~{}\middle|~{}W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 1\right\}$
(4.5) $\mathcal{P}^{4}(x,a):=$$\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~{}\middle|~{}W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 2\right\}$
which are Wasserstein-balls around the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [27, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$.
We depict the trained actions and the associated convergence time (duration of running $1~{}000~{}000$ iterations of the algorithm) in Table 1.
$X_{t}$ 0 1 2 3 4 5 6 7 8 9 10 Convergence time
[…]
Table 1. The trained actions $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$, $a_{t}^{\mathcal{P}^{3}}$, $a_{t}^{\mathcal{P}^{4}}$, $a_{t}^{\rm non-robust}$, in dependence of the realized state $X_{t}$
We next test the profit of the actions described in Table 1 by playing $100~{}000$ rounds of the game according to the trained policies $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$ ,$a_{t}^{\mathcal{P}^{3}}$ and $a_{t}^{\mathcal{P}^{4}}$. For simulating the $100~{}000$ rounds we assume an underlying binomial distribution $\mathbb{P}=\operatorname*{Bin}(10,p_{\rm true})$ with a fixed probability $p_{\rm true}$ for heads which we vary from $0.1$ to $0.9$. We depict the cumulative profits of the considered actions in Table 3.
$p_{\rm true}$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Robust, $\mathcal{P}^{1}$ $-31240$ $-19121$ $-4283$ $15161$ 3048430484{30484}30484 $\mathbf{29857}$ 1321213212{13212}13212 $-10339$ ${-29843}$
Robust, $\mathcal{P}^{2}$ $\mathbf{-24424}$ $\mathbf{5231}$ $\mathbf{21051}$ $\mathbf{24613}$ 2299822998{22998}22998 $11715$ $-4984$ $-18896$ $-31185$
Robust, $\mathcal{P}^{3}$ $-30003$ $-10859$ $9843$ $21900$ $25395$ $22587$ $9653$ $-10656$ $-30301$
Robust, $\mathcal{P}^{4}$ $-24528$ $4263$ $16771$ $13415$ $9925$ $13366$ $\mathbf{17345}$ $\mathbf{4731}$ $\mathbf{-24452}$
Non-Robust $-31101$ $-18251$ $-521$ $23175$ $\mathbf{35244}$ $23227$ $-1387$ $-18421$ $-31024$
Table 2. Overall profit of the game described in Section 4.1 in dependence of different trained strategies (described in Table 1) and of the probability distribution $P=\operatorname*{Bin}(10,p_{\rm true})$ of the simulated underlying process. We indicate with bold characters which is the best performing strategy for each choice of $p_{\rm true}$.
We first observe that only in the case $p_{\rm true}=0.5$ the non-robust strategy is the best performing strategy. The Wasserstein-strategies are outperformed by the robust strategies taking into account either $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$ in all cases for which $p_{\rm true}\leq 0.6$, in particular for the cases $p_{\rm true}=0.3,0.6$ which are by construction contained in the respective ambiguity sets.
The results imply that the choice of the optimal ambiguity sets depends on the scenarios that are deemed possible: If we know that [^5] This would correspond to the case that the dealer in a casino does not tell whether she uses a biased coin (for which we know that $p_{\rm true}=0.6$ for playing the game with an agent or if she uses a fair coin. $p_{\rm true}\in\{0.5,0.6\}$, then one should exactly consider $\mathcal{P}^{1}$, whereas in that situation the set $\mathcal{P}^{3}$ that also covers the case $p_{\rm true}=0.6$ will lead to a suboptimal because too conservative strategy that also covers the case $p_{\rm true}=0.4$ as a possible scenario.
If an applicant however simply has a best guess of a reference measure and considers the cases $p_{\rm true}=0.4$ and $p_{\rm true}=0.6$ both as equally possible, then it is advisable to pursue the Wasserstein-approach outlined in [27].
4.2. Stock investing
This example is a variant of [28, Example 4.3] showcasing the use of tailored ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [6], [22], and [44].
We consider the problem of optimally investing in the stock market, and, to this end, encode the return of a stock by $2$ numeric values: either the return is positive ($1$) or negative ($-1$). The space of numerically encoded returns is therefore given by: $T:=\{-1,1\}$. At each time step $t$ the agent can choose to buy the stock or not, or to short sell the stock. The action $a_{t}$ represent this investment decision, buying the stock is encoded by $1$ and not buying it by 00, whereas short-selling corresponds to $-1$, i.e., we have $A:=\{-1,0,1\}$.
The agent’s investment decision should depend not only on the most recent return but naturally will depend on the current market situation. Therefore the agent relies her investment decision on the last $h=5$ returns. Hence, we consider the state space
$\mathcal{X}:=T^{h}=\{-1,1\}^{h}.$
The reward is given by the function:
(4.6) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a\cdot x^{\prime},$
i.e., we reward a correct investment decision and penalize an incorrect investment decision.
Now, to train the agent, we construct an estimated reference measure in the following way. For the state transition, we consider the historic evolution of the (numerically encoded) returns of the underlying stock. This time series is denoted by $(R_{j})_{j=1,...,N}\subset T^{N}$ for some $N\in\mathbb{N}$. We then define for some small $\gamma>0$ a map [^6] We only employ $\gamma$ to avoid division by 00, hence $\gamma$ can be chosen arbitrarily small. $\mathcal{X}\times A\ni(x,a)\mapsto\widetilde{\mathbb{P}}(x,a):=\sum_{i\in T}p_{i}(x)\cdot\delta_{\{i\}}\in\mathcal{M}_{1}(T)$ where for $x\in\mathcal{X}$, $i\in T$ we define:
(4.7) $p_{i}(x):=\frac{{\widetilde{p}_{i}}(x)+\frac{\gamma}{4}}{\gamma+\sum_{j\in T}\widetilde{p}_{j}(x)}\in[0,1],$
with
(4.8) $\widetilde{p}_{i}(x):=\sum_{j=1}^{X-h+1}1\mathrm{l}_{\{(\pi(x),i)=(R_{j},...,R_{j+h-1})\}},$
and where $\mathbb{R}^{h}\ni(x_{1},\dots,x_{h})\mapsto\pi(x_{1},\dots,x_{h}):=(x_{2},\dots,x_{h})\in\mathbb{R}^{h-1}$ denotes the projection onto the last $h-1$ components. This means the data-driven construction of the measure ~$\widetilde{\mathbb{P}}(x,a)$ relies on the relative frequency of the sequence $(\pi(x),i)$ in the time series of past realized returns $(R_{j})_{j=1,...,N}$ for $i\in\{-1,1,\}$: the more often the market situation $(\pi(x),i)\in\mathcal{X}$ was observed in the training data $(R_{j})_{j=1,\dots,N}$ the higher the value $\widetilde{p}_{i}$. Equation (4.7) is then applied to convert the frequencies $\widetilde{p}_{i}$ to probabilities $p_{i}$. Then, the state transition of the reference measure is partially deterministic ($h-1$ components of subsequent states coincide) and is of the form
(4.9) $X\times A\ni(x,a)\mapsto\mathbb{P}(x,a)=\delta_{\pi(x)}\otimes\widetilde{\mathbb{P}}(x,a)\in\mathcal{M}_{1}(\mathcal{X}).$
To construct the probability measure from (4.9) using real data, we consider daily returns of the stock of Alphabet Inc (previously known as Google) in the time period from beginning of $2010$ until $2024$, where we depict the cumulative sum of the signs of the returns in Figure 1.
Figure 1. Cumulative sum of the sign of the returns and the decomposition of the data in training periods and testing periods.
We split the data into a training period ranging from January $2010$ until December $2021$ and two testing period starting directly thereafter until February $2023$ and January $2024$, respectively. We then first construct a measure $\mathbb{P}_{1}$ according to (4.9) by using all training data. To construct an ambiguity set $\mathcal{P}$ we isolate specific periods of the training data where the market behaved very volatile (Training Period 1), bearish (Training Period 2) or bullish (Training Period 3). Including the accompanying probability measures $\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ in an ambiguity set $\mathcal{P}:=\{\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}\}$ then accounts for the fear of the agent to not trade profitably in these difficult market scenarios. The choice of these specific training periods is depicted in Figure 1. Using Algorithm 1, we then train different agents using the ambiguity set $\mathcal{P}$ and using the non-robust state transitions $\mathbb{P}_{i}$, $i=1,2,3,4$, respectively. Moreover, we consider a trend following strategy (i.e., always invest according to the sign of the last return) as well as a simple buy-and-hold strategy. The results (mean average reward per trade) of the different strategies evaluated in the test period are shown in Table 3, and they show that using a robust strategy indeed enables to avoid losses in both test periods. Note moreover that in the bearish test period 1 the robust strategy in particular outperforms the non-robust agent trained on all data which is certainly due to the inclusion of $\mathbb{P}_{2}$ and $\mathbb{P}_{3}$ in the ambiguity set $\mathcal{P}$ which are the best performing strategies in this period. Also note that in test period 2 it becomes visible that the robust approach due to its construction of a worst-case approach is better suited to avoid losses than to profit from bearish market scenarios.
Test Period 1 Test Period 2
Robust ($\mathcal{P}$) 0.0205 0.0521
Non-Robust: All Data ($\mathbb{P}_{1}$) -0.0274 0.0729
Non-Robust: Training Period 1 ($\mathbb{P}_{2}$) 0.0377 0.0885
Non-Robust: Training Period 2 ($\mathbb{P}_{3}$) 0.0685 0.1042
Non-Robust: Training Period 3 ($\mathbb{P}_{4}$) -0.0034 0.1458
Trend following -0.0137 0.0781
Buy and hold -0.0822 0.1146
Table 3. A comparison of the average reward per trade in the two test periods depicted in Figure 1 among different strategies. The best-performing strategy is highlighted with bold letters.
5. Proofs and auxiliary results
5.1. Auxiliary Results
For any function $f:\mathcal{X}\times A\rightarrow\mathbb{R}$, we write
(5.1) $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}\sup_{a\in A}|f(x,a)|.$
Next, we introduce the operator $\mathcal{H}$ operating on a function $v:\mathcal{X}\times A\rightarrow\mathbb{R}$ being defined by
(5.2) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}v(x,a):=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}v(X_{1},b)\right].$
Lemma 5.1.
Let $10<\alpha<$ 1, and let the ambiguity set $\mathcal{P}$ be defined in (2.1). Then the following fixed-point equation holds true for the optimal $Q$-value function defined in (2.5):
$\mathcal{H}Q^{*}(x,a)=Q^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A.$
Proof.
This follows directly by definition of $Q^{*}$ and by Proposition 2.1, compare also [27, Proof of Lemma 17]. ∎
Lemma 5.2.
For any maps $q_{i}:\mathcal{X}\times A\rightarrow\mathbb{R},\ i=1,2,$ we have:
$\|\mathcal{H}q_{1}-\mathcal{H}q_{2}\|_{\infty}\leq\alpha\|q_{1}-q_{2}\|_{\infty}.$
Proof.
The proof follows is analogue to [27, Proof of Lemma 18]. ∎
The following lemma is a result from stochastic approximation theory adjusted to our setting, compare [27, Lemma A.1] for the presented formulation adjusted to the underlying setting. We refer also to [36, Lemma 1], [14] [18, Theorem 1], [38, Lemma 12], and [42, Lemma 3] for further reference.
Lemma 5.3 ([36], Lemma 1).
Let $\mathbb{P}_{0}\in\mathcal{M}_{1}(\Omega)$ be a probability measure on $(\Omega,\mathcal{F})$, and consider a family of stochastic processes $(\gamma_{t}(x,a),F_{t}(x,a),\Delta_{t}(x,a))_{t\in\mathbb{N}_{0}}$, $(x,a)\in\mathcal{X}\times A$, satisfying for all $t\in\mathbb{N}_{0}$
$\Delta_{t+1}(x,a)=\left(1-\gamma_{t}(x,a)\right){\Delta}_{t}(x,a)+\gamma_{t}(x,a)F_{t}(x,a)\qquad\mathbb{P}_{0}\text{-almost surely for all }(x,a)\in\mathcal{X}\times A.$
Let $(\mathcal{G}_{t})_{t\in\mathbb{N}_{0}}\subseteq\mathcal{F}$ be a sequence of increasing $\sigma$-algebras such that for all $(x,a)\in\mathcal{X}\times A$ the random variables $\Delta_{0}(x,a)$ and $\gamma_{0}(x,a)$ are $\mathcal{G}_{0}$-measurable and such that $\Delta_{t}(x,a)$, $\gamma_{t}(x,a)$, and $F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable for all $t\in\mathbb{N}$. Further assume that the following conditions hold.
  • (i)
    $0\leq\gamma_{t}(x,a)\leq 1$, $\sum_{t=0}^{\infty}\gamma_{t}(x,a)=\infty$, $\sum_{t=0}^{\infty}\gamma_{t}^{2}(x,a)<\infty$ $\mathbb{P}_{0}$-almost surely for all $(x,a)\in\mathcal{X}\times A$, $t\in\mathbb{N}_{0}$.
  • (ii)
    There exists $\delta\in(0,1)$ such that $\|\mathbb{E}_{\mathbb{P}_{0}}\left[F_{t}(\cdot,\ \cdot)~{}\middle|~{}\mathcal{G}_{t}\right]\|_{\infty}\leq\delta\|\Delta_{t}\|_{\infty}$ $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$.
  • (iii)
    There exists $C>0$ such that $\left\|\operatorname{Var}_{\mathbb{P}_{0}}\left(F_{t}(\cdot,\ \cdot)~{}\middle|~{}\mathcal{G}_{t}\right)\right\|_{\infty}\leq C(1+\|\Delta_{t}\|_{\infty})^{2}$ $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$.
Then, $\lim_{t\rightarrow\infty}\Delta_{t}(x,a)=0\qquad\mathbb{P}_{0}$-almost surely for all $(x,a)\in\mathcal{X}\times A$.
The following result is usually referred to as Popoviciu’s inequality, see [31] or [33].
Lemma 5.4.
Let $Z$ be a random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ satisfying $m\leq Z\leq M$ for some $-\infty<m\leq M<\infty$. Then, we have
$\operatorname{Var}_{\mathbb{P}}(Z)\leq\frac{1}{4}\left(M-m\right)^{2}.$
5.2. Proofs of the results from Section 2 and Section 3
In this section we provide the proofs of Section 2.4 and Section 3.
Proof of Proposition 2.1.
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [28, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [28, Assumption 2.2] and [28, Assumption 2.4] by showing the following four properties.
  • (i)
    The set-valued map
    $\mathcal{X}\times A$ $\rightarrow(\mathcal{M}_{1}(\mathcal{X})^{N},\tau_{0})$
    $(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a):=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\}$
    is nonempty, compact-valued and continuous [^7] Continuity of a set-valued map means the map is both upper hemicontinuous and lower hemicontinuous..
  • (ii)
    There exists $C_{P}\geq 1$ such that for all $(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)$ we have
    (5.3) $\int_{X}(1+\|y\|^{p})\mathbb{P}(dy)\leq C_{P}(1+\|x\|^{p})$
  • (iii)
    The reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$ is Lipschitz-continuous in its first two components.
  • (iv)
    We have $0<\alpha<\frac{1}{C_{P}}$, where $C_{P}$ is defined in (ii).
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty and since the set $\mathcal{P}(x,a)$ is finite it is also compact for all $(x,a)\in\mathcal{X}\times A$. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~{}|~{}(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both, $\mathcal{X}$ and $A$ are finite, we have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
To show the lower hemicontinuity let $(x,a)\in\mathcal{X}\times A$, $\mathbb{P}\in\mathcal{P}(x,a)$ and consider a sequence $(x_{n},a_{n})_{n\in\mathbb{N}}\subseteq\mathcal{X}\times A$ with $(x_{n},a_{n})\rightarrow(x,a)$ as $n\rightarrow\infty$. By definition of $\mathcal{P}$, there exists some $i\in\{1,\dots,N\}$ such that $\mathbb{P}^{(i)}(x,a)=\mathbb{P}$. We use this observation to define
$\mathbb{P}_{n}:=\mathbb{P}^{(i)}(x_{n},a_{n})\in\mathcal{P}(x_{n},a_{n}),~{}n\in\mathbb{N},$
and we obtain, since $(x_{n},a_{n})=(x,a)$ for $n$ large enough that $\mathbb{P}_{n}\rightarrow\mathbb{P}$ weakly which shows the lower hemicontinuity of $\mathcal{P}$ according to [2, Theorem 17.21].
To show (ii), note that $p=0$, and define $C_{P}:=1$, then we have
$\int_{\mathcal{X}}(1+\|y\|^{0})\mathbb{P}(dy)=2=C_{P}\cdot(1+\|x\|^{0})$
for all $(x,a)\in\mathcal{X}\times A$. Next, (iii) follows since for all $x_{0},x_{0}^{\prime},x_{1}\in\mathcal{X}$ and $a,a^{\prime}\in A$ we have
$|r(x_{0},a,x_{1})-r(x_{0}^{\prime},a^{\prime},x_{1})|\leq\left(\max_{y_{0},y_{0}^{\prime}\in\mathcal{X},~{}b,b^{\prime}\in A\atop(y_{0},b)\neq(y_{0}^{\prime},b^{\prime})}\frac{|r(y_{0},b,x_{1})-r(y_{0}^{\prime},b^{\prime},x_{1})|}{\|y_{0}-y_{0}^{\prime}\|+\|b-b^{\prime}\|}\right)\cdot\left(\|x_{0}-x_{0}^{\prime}\|+\|a-a^{\prime}\|\right).$
Finally, (iv) follows since $C_{P}=1$ and as we have $10<\alpha<$ 1 by assumption. ∎
Proof of Theorem 3.1.
To establish convergence of the $Q$-learning algorithm from Algorithm 1, we aim at applying Lemma 5.3. To verify the assumptions of Lemma 5.3 we first note that for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ the update rule from (3.2) can be written equivalently as
(5.4) $Q_{t+1}(x,a)=Q_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big{)};$
where for $(x,a)\in\mathcal{X}\times A$ we recall the definition
(5.5) $\gamma_{t}(x,a,X_{t}):=1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{t}.$
Next, for for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ we define
$\Delta_{t}(x,a):$ $=Q_{t}(x,a)-Q^{*}(x,a),$
$F_{t}(x,a):$ $=\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)1\mathrm{l}_{\left\{(x,a)=\left(X_{t},~{}a_{t}(X_{t})\right)\right\}}.$
Moreover, we recall that as defined in Algorithm 1 we have
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]$
Once we have verified that the three assumptions of Lemma 5.3 are fulfilled, we obtain $\Delta_{t}(x,a)\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely as $t\rightarrow\infty$ and hence
$Q_{t}(x,a)\rightarrow Q(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
implying the assertion. Thus, to apply Lemma 5.3, we compute for all $(x,a)\in\mathcal{X}\times A$ that
$\Delta_{t+1}(x,a)=$ $Q_{t}(x,a)-Q^{*}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big{)}$
$=$ $\Delta_{t}(x,a)$
$+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)+Q^{*}(x,a)-Q^{*}(x,a)\Big{)}$
$=$ $(1-\gamma_{t}(x,a,X_{t}))\cdot\Delta_{t}(x,a)$
$+\gamma_{t}(x,a,X_{t})\cdot\Big{(}r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\Big{)}$
$=$ $(1-\gamma_{t}(x,a,X_{t}))\cdot\Delta_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot F_{t}(x,a),$
where we used in the last step that the indicator $1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}$ appears both in the definition of $F_{t}$ and $\gamma_{t}$ and that the square of the indicator function is the indicator function itself. Consider the filtration $(\mathcal{G}_{t})_{t\in\mathbb{N}}$ with
$\mathcal{G}_{t}:=\sigma({X_{1},X_{2},...,X_{t}}),\ t\in\mathbb{N},$
and $\mathcal{G}_{0}:=\{\varnothing,\Omega^{N}\}$ being the trivial sigma-algebra. Then, by definition we have for all $t\in\mathbb{N}$ and for all $(x,a)\in\mathcal{X}\times A$ that the random variables $\Delta_{t}(x,a),\ \gamma_{t}(x,a),\ F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable. Next, let $(x,a)\in\mathcal{X}\times A,$ and $t\in\mathbb{N}$. Then, we have by the definition of $\mathbb{P}_{x_{0},\textbf{a}}$ in (3.3) $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely that
$|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)~{}\middle|~{}\mathcal{G}_{t}\right]\right|$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right]\right|$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\Bigg{|}\min_{i=1,...,N}\left(\mathbb{E}_{\mathbb{P}^{(i)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right)-Q^{*}(x,a)\Bigg{|}$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\Bigg{|}\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q_{t}(X_{1},b)\right]-Q^{*}(x,a)\Bigg{|}.$
By using the definition of the operator $\mathcal{H}$ from (5.2) we therefore obtain
$|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}|\mathcal{H}Q_{t}(x,a)-Q^{*}(x,a)|.$
Consequently, applying Lemma 5.1 we have
(5.6) $|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}|\mathcal{H}Q_{t}(x,a)-\mathcal{H}Q^{*}(x,a)|.$
Now using (5.6) and applying Lemma 5.2 we can conclude
$\|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(\cdot,\cdot)|\mathcal{G}_{t}]\|_{\infty}\leq\|\mathcal{H}Q_{t}-\mathcal{H}Q^{*}\|_{\infty}\leq\alpha\|Q_{t}-Q^{*}\|_{\infty}=\alpha\|\Delta_{t}\|_{\infty}$
showing Lemma 5.3 (ii).
Next, to show that the assumption from Lemma 5.3 (iii) is fulfilled, we observe that for all $(x,a)\in\mathcal{X}\times A$ and $t\in\mathbb{N}$ we have
(5.7) $\operatorname{Var}$ (${}_{\mathbb{P}_{x_{0},\textbf{a}}}(F_{t}(x,a)|\mathcal{G}_{t})$
$=1\mathrm{l}_{\{(x,a)=(X_{t},~{}a_{t}(X_{t}))\}}\cdot\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)$
$\leq\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q^{*}(x,a)\right)\right)$
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right)$
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})\right).$
We define $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and compute as an upper bound for the integrand from (5.7) that
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\leq C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\max_{b^{\prime}\in A}Q^{*}(X_{t+1},b^{\prime})$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}(\max_{b^{\prime}\in A}Q^{*}(y,b^{\prime})-\max_{b\in A}Q_{t}(y,b))$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}\max_{b\in A}|Q^{*}(y,b)-Q_{t}(y,b)|$
$\leq C_{r}+\alpha\|\Delta_{t}\|_{\infty}=:M\in\mathbb{R}.$
On the other hand we also may compute a lower bound as follows
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}+\alpha(\min_{y^{\prime}\in\mathcal{X}}\min_{b\in A}(Q^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)))$
$\geq-C_{r}+\alpha(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)|)$
$\geq-C_{r}+\alpha\min_{y\in\mathcal{X}}(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q^{*}(y^{\prime},b)-Q_{t}(y,b)|)$
$\geq-C_{r}-\alpha(\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q_{t}(y,b)-Q^{*}(y,b)|+|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|)$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|=:m\in\mathbb{R}.$
We apply Popoviciu’s inequality on variances from Lemma 5.4 with the bounds $m$ and $M$ computed just above, to (5.7) and obtain
$\operatorname{Var}_{\mathbb{P}_{x_{0},\textbf{a}}}$ $(F_{t}(x,a)|\mathcal{G}_{t})$
$\leq$ $\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q^{*}(y^{\prime},b^{\prime})\right)$
$\leq$ $\frac{1}{4}(M-m)^{2}$
$\leq$ $\frac{1}{4}\Big{(}C_{r}+\alpha\|\Delta_{t}\|_{\infty}+C_{r}+\alpha\|\Delta_{t}\|_{\infty}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}$
$=$ $\frac{1}{4}\Big{(}C_{r}+C_{r}+2\alpha\|\Delta_{t}\|_{\infty}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}$
$\leq$ $\frac{1}{2}\left(\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\left(\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\Bigg{(}\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}$
$+4\alpha^{2}+\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}\|\Delta_{t}\|_{\infty}^{2}\Bigg{)}$
$\leq$ $\left(4\alpha^{2}+\Big{(}2C_{r}+\alpha\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|\Big{)}^{2}\right)(1+\|\Delta_{t}\|_{\infty}^{2})$
$\leq$ $C\cdot(1+\|\Delta_{t}\|_{\infty})^{2},$
with
(5.8) $C:=\left(4\alpha^{2}+(2C_{r}+\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q^{*}(y,b)-Q^{*}(y^{\prime},b)|)^{2}\right).$
This shows that also Lemma 5.3 (iii) is fulfilled, and hence, with an application of Lemma 5.3 we obtain $Q_{t}(x,a)\rightarrow Q(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$ for all $(x,a)\in\mathcal{X}\times A$. ∎