### v3

# Q-Learning under Finite Model Uncertainty

Abstract.
We propose a robust Q-learning algorithm for Markov decision processes under model uncertainty when each state–action pair is associated with a finite ambiguity set of candidate transition kernels. This finite-measure framework enables highly flexible, user-designed uncertainty models and goes beyond the common KL/Wasserstein-ball formulations. We establish almost sure convergence of the learned Q-function to the robust optimum, and derive non-asymptotic high-probability error bounds that separate stochastic approximation error from transition-kernel estimation error. Finally, we show that Wasserstein-ball and parametric ambiguity sets can be approximated by finite ambiguity sets, allowing our algorithm to be used as a generic solver beyond the finite setting.
Keywords: Q-Learning, Markov Decision Processes, Model Uncertainty
1National University of Singapore, Department of Mathematics,
21 Lower Kent Ridge Road, 119077.
Corresponding author, email:jul_\_ses@nus.edu.sg
1. Introduction
Markov decision processes (MDPs) are a fundamental framework for sequential decision making under uncertainty, and, in particular, for reinforcement learning. An agent repeatedly observes the state of a stochastic environment, selects an action, and receives a reward, with the goal of maximizing cumulative discounted rewards over an infinite time horizon. Central to this formulation is the specification of the state transition probabilities, which determine how actions influence the evolution of the system.
In practice, these transition probabilities are rarely known and must be estimated from data ([1], [44], [50]) or implicitly learned through interaction with the environment (see, e.g., [10], [53]). As a consequence, classical (non-robust) reinforcement learning methods are highly sensitive to model misspecification: policies trained under an incorrect or outdated transition model may perform poorly when deployed, even if the deviation from the training model is moderate. This issue is particularly acute in non-stationary environments or in applications where data are scarce and multiple competing models are plausible. A natural way to address this challenge is to incorporate model uncertainty directly into the training phase. In distributionally robust Markov decision problems ([38]), the transition kernel is assumed to belong to an ambiguity set of probability measures, and the agent optimizes performance under the worst-case model within this set. This paradigm reflects Knightian uncertainty ([26]) and provides policies that are robust against adverse but plausible deviations of the true dynamics.
Most existing robust reinforcement learning approaches specify ambiguity sets as metric balls, such as Kullback–Leibler ([7], [29]) or Wasserstein balls ([39]), centered at a reference transition kernel. While mathematically convenient, this construction implicitly enforces a neighborhood-based and typically symmetric notion of uncertainty. As a result, it does not allow the decision maker to precisely control which alternative transition models are considered plausible, nor to express asymmetric or scenario-based uncertainty.
In many applications, however, uncertainty is better described by a finite collection of competing models. These may arise from different estimation procedures, stress scenarios, regime-specific dynamics, or expert judgments. In such settings, it is often desirable to protect explicitly against a small number of adverse scenarios rather than against all distributions within a metric neighborhood of a reference model. This motivates the study of robust control problems with finite, user-specified ambiguity sets.
To illustrate the systematic difference, consider an agent repeatedly betting on the outcome of a coin toss game, where the state records recent outcomes and the agent chooses whether to bet on an increase or decrease of the next realization. Suppose the true data-generating process is unknown but believed to belong to a small set of plausible models, such as a fair or a moderately biased coin. A non-robust agent trained under a single reference model may perform well when this model is correct, but can suffer systematic losses under alternatives. Wasserstein-robust approaches mitigate this by protecting against all distributions in a neighborhood of the reference model; however, this typically introduces excessive conservatism by including many unintended intermediate distributions. In contrast, finite ambiguity sets allow the decision maker to specify explicitly the competing scenarios and to optimize against their worst case. As shown in Section 6, this scenario-based robustness yields qualitatively different policies and can outperform both non-robust and Wasserstein-robust strategies when uncertainty is concentrated on a small number of concrete models.
In this paper, we develop a robust $Q$-learning framework for Markov decision problems in which each state–action pair is associated with a finite set of candidate transition kernels. We allow the transition kernel selected by nature to vary over time, as long as it remains within the prescribed ambiguity set, leading to a time-inhomogeneous worst-case formulation. This setting preserves the dynamic programming principle while enabling highly flexible modeling of uncertainty.
1.1. Our contribution
Our main contributions can be summarized as follows.
  1. 1.)
    We introduce a distributionally robust MDP framework based on finite ambiguity sets of transition probabilities. This approach enables scenario-based robustness and allows the decision maker to specify explicitly which alternative models are deemed plausible, without being restricted to metric balls around a reference measure.
  2. 2.)
    We propose a novel robust $Q$-learning algorithm for this setting. The algorithm requires only generative access to each candidate transition kernel and does not assume explicit knowledge of the underlying distributions. At each iteration, the learner evaluates actions under the empirically worst-case model and updates the Q-function accordingly.
  3. 3.)
    We provide a rigorous theoretical analysis of the proposed algorithm. We establish almost sure convergence of the learned Q-function to the robust optimal Q-function and derive non-asymptotic high-probability error bounds. These bounds naturally decompose into a stochastic approximation error and an additional term reflecting the statistical error arising from estimating the candidate transition kernels. In the special case where the kernels are known, this latter term vanishes.
  4. 4.)
    We show that a broad class of ambiguity sets, including Wasserstein balls and parametric families of transition kernels, can be approximated by sequences of finite ambiguity sets. These results allow the proposed algorithm to serve as a generic numerical solver beyond the finite setting.
  5. 5.)
    We show tractability of our approach through two examples: one coin toss example allowing to fully understand the presented approach and how it distinguishes from other related approaches as well as a stock allocation example using real data.
1.2. Related work
Distributionally robust optimization and robust Markov decision problems have been studied extensively over the past decade (compare, e.g., [7], [8], [17], [31], [38], [41], [42], [47], [48], [55], [60], [64], [65], [66], and [67]). This literature has primarily focused on establishing dynamic programming principles under model uncertainty, characterizing robust value functions, and developing tractable solution methods for specific classes of ambiguity sets. A central modelling choice in these works is the specification of uncertainty via ambiguity sets of transition kernels, typically designed to capture deviations from a reference model in a principled manner.
Classical Q-learning for Markov decision problems was introduced by Watkins ([62] [61]), and has since become one of the cornerstone algorithms in reinforcement learning. Comprehensive introductions can be found in [13] and [24], and a wide range of applications has been studied, including control, finance, and operations research (see, e.g., [3], [4], [11], [12], [21], [27], and [37]). These approaches, however, assume that the underlying transition kernel is fixed and correctly specified, making them sensitive to model misspecification.
Only more recently has attention shifted to robust Q-learning methods that explicitly account for uncertainty in the transition dynamics. Existing approaches predominantly consider ambiguity sets defined as metric balls—most notably Kullback–Leibler balls ([8], [29], and [59]) and Wasserstein balls ([38])—around a reference transition kernel. These formulations exploit convex duality and the structure of the underlying metric to obtain tractable algorithms and theoretical guarantees. At the same time, metric-ball ambiguity sets impose a neighbourhood-based and typically symmetric notion of uncertainty, which implicitly includes many transition models that may not be considered plausible by a decision maker.
In contrast, many practical applications call for scenario-based uncertainty modelling, where ambiguity is naturally described by a finite collection of competing transition models arising from different estimators, stress scenarios, regimes, or expert assessments. Finite ambiguity sets are also particularly natural in financial applications, where uncertainty is often driven by a small number of stress scenarios or regime-dependent models rather than diffuse perturbations of a reference measure, see, e.g. [20], [25], and [45].
To the best of our knowledge, robust $Q$-learning methods that allow for arbitrary finite collections of candidate transition kernels, without relying on metric-ball constructions or convexity assumptions, have not been developed. The present work fills this gap by providing an algorithmic and theoretical treatment of robust $Q$-learning under finite model uncertainty, thereby enabling a flexible, user-designed approach to robustness that complements existing metric-based methods.
1.3. Structure
The remainder of the paper is organized as follows. Section 2 introduces the robust Markov decision problem with finite ambiguity sets and establishes the associated dynamic programming formulation. In Section 3, we present the robust $Q$-learning algorithm and prove its almost sure convergence. Section 4 shows how ambiguity sets consisting of infinitely many probability measures can be approximated by finite sets, including Wasserstein and parametric ambiguity models. Section 5 outlines an extension to continuous state spaces using function approximation and deep $Q$-learning. Section 6 provides numerical experiments illustrating the tractability and qualitative behavior of the proposed approach. All proofs are collected in Section 7.
2. Setting and specification of the problem
In this section we introduce the underlying framework which will be employed to establish a robust $Q$-learning algorithm for finite ambiguity sets.
2.1. Setting
Optimal control problems, such as Markov decision problems, are formulated using a state space comprising all possible states accessible by an underlying stochastic process. We model this state space by a $d$-dimensional finite Euclidean subset $\mathcal{X}\subset\mathbb{R}^{d}$ and eventually aim at solving a robust control problem over an infinite time horizon. Hence, we define the state space over the entire time horizon via
$\Omega:=\mathcal{X}^{\mathbb{N}}=\mathcal{X}\times\mathcal{X}\times$...
equipped with the corresponding $\sigma$-algebra $\mathcal{F}:=2^{\mathcal{X}}\otimes 2^{\mathcal{X}}\otimes\cdots$. Next, let $(X_{t})_{t\in\mathbb{N}}$ be the state process, i.e., the stochastic process on $\Omega$ describing the states attained over time. We denote the finite set of possible actions by $A\subset\mathbb{R}^{m}$, where $m\in\mathbb{N}$ represents the dimension of the action space. The set of admissible policies is then given by
$\mathcal{A}:=$ $\{\textbf{a}=(a_{t})_{t\in\mathbb{N}}\ |\ (a_{t})_{t\in\mathbb{N}}:\Omega\rightarrow A;\ a_{t}\text{ is }\ \sigma(X_{t})-\text{measurable}\ \text{for all }t\in\mathbb{N}\}$
$=$ $\{(a_{t}(X_{t}))_{t\in\mathbb{N}}\ |\ a_{t}:\mathcal{X}\rightarrow A\ \text{Borel\ measurable }\text{for all }t\in\mathbb{N}\}.$
In contrast to classical non-robust Markov decision problems, we work under the paradigm that the precise state transition probability is not known but instead, to account for model uncertainty, contained in an ambiguity set of finitely many transition probabilities. In the following, let $\mathcal{M}_{1}(\mathcal{X})$ denote the set of probability measures on $(\mathcal{X},\mathcal{F})$, and let $\tau_{0}$ denote the topology of weak convergence [^1] For any $\mu\in\mathcal{M}_{1}(X)$, $(\mu_{n})_{n\in\mathbb{N}}\subseteq\mathcal{M}_{1}(X)$ and $C_{b}(X,\mathbb{R})$ denoting the space of continuous and bounded functions mapping from the space $X$ to $\mathbb{R}$, we have: $\underset{n\rightarrow\infty}{\mu_{n}\longrightarrow^{\tau_{0}}\mu}\Leftrightarrow\lim\limits_{n\to\infty}\int gd\mu_{n}=\int gd\mu,\ \text{ for all }g\in C_{b}(X,\mathbb{R}).$ . An ambiguity set of $N\in\mathbb{N}$ probability measures in dependence of a state-action pair is then modeled by a set-valued map given by
(2.1) $\mathcal{X}\times A$ $\rightarrow(\mathcal{M}_{1}(\mathcal{X})^{N},\tau_{0})$
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\},$
where for all $k\in\{1,\dots,N\}$ and for all $(x,a)\in\mathcal{X}\times A$ we have that $\mathbb{P}^{(k)}(x,a)$ is a probability measure, i.e., $\mathbb{P}^{(k)}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. Note that in the degenerate case where $\mathcal{P}(x,a)$ contains only a single probability distribution, we are facing a classical Markov decision process (i.e. a non-robust one), compare, e.g., [9].
The ambiguity set of admissible probability distributions on $\Omega$ depends on the initial state $x\in\mathcal{X}$ and the policy $\textbf{a}\in\mathcal{A}$. Let $\delta_{x}\in\mathcal{M}_{1}(\mathcal{X})$ denote the Dirac measure at point $x\in\mathcal{X}$. Then, we define for every state-action pair $(x,\textbf{a})\in\mathcal{X}\times\mathcal{A}$ the underlying set of admissible probability measures of the stochastic process $(X_{t})_{t\in\mathbb{N}}$ by
[…]
$\text{and }\mathbb{P}_{t}\in\mathcal{P}(x_{t},a_{t}(x_{t})),\ \text{ for all }x_{t}\in\mathcal{X}\Bigl\},$
where the notation $\mathbb{P}=\delta_{x}\otimes\mathbb{P}_{0}\otimes\mathbb{P}_{1}\otimes...\in\mathfrak{P}_{x,a}$ abbreviates the infinite concatenation of the conditional probability distributions:
$\mathbb{P}(B)=\sum_{x_{0}\in\mathcal{X}}...\sum_{x_{t}\in\mathcal{X}}...1\hskip-3.0pt\mathrm{l}_{B}((x_{t})_{t\in\mathbb{N}})...\mathbb{P}_{t-1}(x_{t-1};\{x_{t}\})...\mathbb{P}_{0}(x_{0};\{x_{1}\})\delta_{x}(\{x_{0}\}),\qquad B\in\mathcal{F}.$
The construction of ambiguity sets $\mathfrak{P}_{x,\textbf{a}}$ provided in (2.1) and (2.2) is in the related literature often referred to as a $(s,a)$-rectangular ambiguity set, (compare [22] and [64]) and enables to obtain a dynamic programming principle of the associated Markov decision problem which allows to derive tractable numerical solution methods, compare Section 2.4.
2.2. Finite ambiguity sets and other ambiguity sets
Note that the paradigm under which we work in this paper is fundamentally different from the approaches pursued, e.g., in [8], [29], and [39] where the corresponding ambiguity sets of probability measures are defined as balls around some reference measure.
The $Q$-learning approaches presented in [8], [29], and [39] provide solutions to account for a potential misspecification of a reference measure, by allowing deviations from it where the size of the deviations is measured by the respective distances (Kullback–Leibler distance and Wasserstein distance). However, these approaches allow not to control of which type the distributions in the ambiguity set are but instead consider all measures that are in a certain sense close to the reference measure. If a decision maker is instead interested in controlling the distributions contained in the ambiguity set, for example, by allowing only for a specific type of parametric distributions with a finite set of possible parameters or by considering an asymmetric ambiguity set, the decision maker can use the $Q$-learning approach presented in this paper. Another relevant situation is the consideration of multiple estimated models / probability measures obtained for exampled via different estimation methods. Wishing to take into account all of these estimations, a decision maker can construct an ambiguity set containing all estimated probability measures.
Beyond these use cases, our approach also allows to combine different types of distributions, if desired.
The main difference therefore is that a decision maker using the approach presented in this paper can precisely control what types of distributions are deemed possible whereas the approaches from [8], [29], and [39] account for a misspecification of one single reference measure.
2.3. Specification of the optimization problem
After execution of an action $a_{t}(X_{t})$, the agent receives a feedback on the quality of the chosen action in terms of a reward $r(X_{t},a_{t}(X_{t}),X_{t+1})$ where $r$ denotes some reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$. The robust optimization problem consists then, for every initial state $x\in\mathcal{X}$, in maximizing the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ over all possible policies $\textbf{a}\in\mathcal{A}$, where $\alpha\in(0,1)$ is some discount factor accounting for time preferences of rewards. Therefore, the value function
(2.3) $\mathcal{X}\ni x\mapsto V(x):=\sup_{\mathbf{a}\in\mathcal{A}}\inf_{\mathbb{P}\in\mathfrak{P}_{x,\textbf{a}}}\mathbb{E}_{\mathbb{P}}\left[\sum^{\infty}_{t=0}\alpha^{t}\cdot r(X_{t},a_{t}(X_{t}),X_{t+1})\right]$
describes the expected value of $\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},a_{t}(X_{t}),X_{t+1})$ under the worst case measure from $\mathfrak{P}_{x,\textbf{a}}$ and when executing the optimal policy $\textbf{a}\in\mathcal{A}$ after having started in initial state $x\in\mathcal{X}$.
2.4. Dynamic programming
To solve (2.3) directly is typically a non-feasible optimization problem as it means to find directly (infinite-dimensional) solutions over the whole time horizon. However, due to its time-homogeneous structure, solving the infinite time horizon problem can be simplified significantly to a one time step problem which turns out to be tractable. To this end, we consider the one time step optimization problem
(2.4) $\mathcal{X}\ni x\mapsto\mathcal{T}V(x):=\max_{a\in A}\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
where $\mathcal{X}\ni x\mapsto V(x)$ is the value function defined in (2.3). By using the above definitions one can show (see Proposition 2.1) that the dynamic programming equation $\mathcal{T}V=V$ holds, implying that solving the fixed point equation $\mathcal{T}V=V$ is sufficient to compute the optimal value function. This fixed-point equation is also famously known in its non-robust formulation as Bellman equation, compare, e.g. [14, p.164] for a non-robust version.
To determine optimal actions, analogous to the definition of the non-robust $Q$-value function (see e.g. [62]), we further define the robust optimal $Q$-value function by
(2.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q^{*}(x,a):=\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
allowing us to interpret $Q^{*}(x,a)$ as the quality of executing action $a$ when in state $x$ (under the worst case measure) with the best possible action leading to $V(x)$ explaining the notion of quality.
Proposition 2.1.
Assume $0<\alpha<1$. Then, for all $x\in\mathcal{X}$ we have
$\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)=V(x).$
Using Proposition 2.1, the goal of our $Q$-learning algorithm presented in the next section will be to algorithmically determine $Q^{*}$ which in turn allows to infer optimal actions in every given state.
3. Robust Q-Learning Algorithm
In this section, we introduce Algorithm 1, a novel distributionally robust $Q$-learning algorithm. We establish sufficient conditions for its convergence in Theorem 3.1 and derive explicit convergence rates in Subsection 3.2.
3.1. The robust Q-Learning algorithm and its convergence
We do not assume explicit knowledge of the candidate transition kernels in $\mathcal{P}$. Instead, we only require generative access: for every $(x,a)\in\mathcal{X}\times A$ and each $k\in\{1,\dots,N\}$, the learner can generate independent samples from the transition kernel $\mathbb{P}^{(k)}(x,a)$. Thus, the algorithm only requires access to $N$ independent simulators, a standard assumption in generative-model reinforcement learning.
Algorithm 1 Robust $Q$-learning for finite ambiguity sets
Input:  State space $\mathcal{X}\subseteq\mathbb{R}^{d}$;  Initial state $x_{0}\in\mathcal{X}$;  Action space $A\subseteq\mathbb{R}^{m}$;  Reward function $r$;  Discount factor $\alpha\in(0,1)$;  Ambiguity set $\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\ldots,\mathbb{P}^{(N)}(x,a)\right\}$;  Policy $(a_{t})_{t}\in\mathcal{A}$;  Sequence of learning rates $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$;
 Initialize $Q_{0}(x,a)$ for all $(x,a)\in\mathcal{X}\times A$ to an arbitrary real value;
 Set $X_{0}=x_{0}$;
 Initialize $\mathbb{P}_{0}^{(k)}(x,a)(y)=\frac{1}{|\mathcal{X}|}$ for all $(x,a,y)\in\mathcal{X}\times A\times\mathcal{X}$;
 Initialize $N_{0}(x,a)=0$ for all $(x,a)\in\mathcal{X}\times A$;
for $t=0,1,\ldots$ do
  Determine an index $k_{t}^{*}\in\{1,\ldots,N\}$ by
(3.1) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\left\{\sum_{y\in\mathcal{X}}\left(r(X_{t},a_{t}(X_{t}),y)+\alpha\max_{b\in A}Q_{t}(y,b)\right)\mathbb{P}_{t}^{(k)}(X_{t},a_{t}(X_{t}))(y)\right\};$
  Sample $X_{t+1}^{(k)}\sim\mathbb{P}^{(k)}\left(X_{t},a_{t}(X_{t})\right)$ independently for $k=1,\dots,N$;
  Set $X_{t+1}:=X_{t+1}^{(k_{t}^{*})}$;
  Set
(3.2) $\mathbb{P}_{t+1}^{(k)}(x,a)(y):=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}\text{ for all }k\in\{1,\dots,N\},~~(x,a,y)\in\mathcal{X}\times A\times\mathcal{X};$
  for all $(x,a)\in\mathcal{X}\times A$ do
   Update $Q_{t}$ via
(3.3) $Q_{t+1}(x,a)\leftarrow\begin{cases}(1-\widetilde{\gamma}_{N_{t}(x,a)})\cdot Q_{t}(x,a)+\widetilde{\gamma}_{N_{t}(x,a)}\cdot\left(r(x,a,X_{t+1})+\alpha\max\limits_{b\in A}Q_{t}(X_{t+1},b)\right)&\text{if }(x,a)=\left(X_{t},a_{t}(X_{t})\right),\\ Q_{t}(x,a);&\text{else};\end{cases}$
and update the number of visit times $N_{t}$ via
$N_{t+1}(x,a)\leftarrow\left\{\begin{array}[]{ll}N_{t}(x,a)+1&\text{if}\ (x,a)=\left(X_{t},~a_{t}(X_{t})\right),\\ N_{t}(x,a)&\text{else};\end{array}\right$.
  end for
end for
Output: A sequence $(Q_{t}(x,a))_{t\in\mathbb{N},x\in\mathcal{X},a\in A}$;
Given a policy $\textbf{a}\in\mathcal{A}$ and some initial state $x_{0}\in\mathcal{X}$ we further define a probability measure by
(3.4) $\mathbb{P}_{x_{0},\textbf{a}}:=\delta_{x_{0}}\otimes\mathbb{P}^{(k_{0}^{*})}(\cdot,a_{0}(\cdot))\otimes\mathbb{P}^{(k_{1}^{*})}(\cdot,a_{1}(\cdot))\otimes\cdots\in\mathfrak{P}_{x_{0},\textbf{a}},$
where $k_{t}^{*}$ is defined in (3.1) and denotes the index of the transition kernel that attains the worst-case expected continuation value (based on the empirical estimate) at time $t$. The measure $\mathbb{P}_{x_{0},\mathbf{a}}$ thus characterizes the (unknown) law of the state process generated by Algorithm 1 under policy $\mathbf{a}$.
Our main result establishes that the sequence $(Q_{t})_{t\in\mathbb{N}}$ produced by Algorithm 1 converges pointwise (in $(x,a)$), $\mathbb{P}_{x_{0},\mathbf{a}}$-almost surely, to the optimal robust $Q$-value function $Q^{*}$ defined in (2.5).
Theorem 3.1.
Let $(\widetilde{\gamma_{t}})_{t\in\mathbb{N}}\subseteq[0,1]$, and define for all $(x,a)\in\mathcal{X}\times A$ and for all $t\in\mathbb{N}$
(3.5) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}.$
where
$N_{t}(x,a):=\sum_{s=0}^{t}\mathbf{1}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}$
is the number of visits to $(x,a)\in\mathcal{X}\times A$ up to time $t$. Moreover, let $0<\alpha<1$ and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(3.6) $\sum^{\infty}_{t=0}\gamma_{t}(x,a,X_{t})=\infty,\ \sum^{\infty}_{t=0}\gamma_{t}^{2}(x,a,X_{t})<\infty,\ \text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Then, we have  $\text{ for all }(x,a)\in\mathcal{X}\times A$ that
(3.7) $\lim_{t\rightarrow\infty}Q_{t}(x,a)=Q^{*}(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Remark 3.2.
If one has explicit knowledge of the distribution $\mathbb{P}^{(k)}$ (and not only a ready-to-use simulator that can sample from the distribution) for each candidate model $k=1,\dots,N$, then (3.1) can be replaced by
(3.8) $k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,\ldots,N}\left\{\sum_{y\in\mathcal{X}}\left(r(X_{t},a_{t}(X_{t}),y)+\alpha\max_{b\in A}Q_{t}(y,b)\right)\mathbb{P}{(k)}(X_{t},a_{t}(X_{t}))(y)\right\};$
and there is no need to learn $\mathbb{P}_{t}$ via (3.2).
Remark 3.3 (On the implementation).
  • (a)
    The idea behind replacing $\widetilde{\gamma}_{t}$ with $\widetilde{\gamma}_{N_{t}(x,a)}$ is to account for the exploration-exploitation trade off by reducing the learning rate the more often we have visited a state and hence the more confident the algorithm is w.r.t.  choosing the correct optimal action, and conversely to apply a larger learning rate if the state-action pair was visited less often until time $t$. This effectively enhances the convergence speed, compare, e.g., [32].
  • (b)
    Note that applying Algorithm 1 requires to choose a pre-determined policy $\textbf{a}\in\mathcal{A}$. Implied from usual practice in non-robust $Q$-learning (([15], Chapter 9), [36], or [54]) and robust $Q$-learning [39], a reasonable choice seems to be the $\varepsilon$-greedy policy defined by
    $\mathcal{X}\ni x\mapsto a_{t}(x):=\begin{cases}\operatorname{argmax}_{b\in A}Q_{t}(x,b)&\text{with probability }1-\epsilon_{\operatorname{greedy}},\\ a\sim\mathcal{U}(A)&\text{with probability }\epsilon_{\operatorname{greedy}},\end{cases}$
    for some $\epsilon_{\operatorname{greedy}}>0$, and where $a\sim\mathcal{U}(A)$ means that a random action $a$ is chosen uniformly at random from the finite set $A$.
3.2. Convergence rates
Analogue to results in classical $Q$-learning ([18], [52]) and in addition to the almost sure convergence guarantee from Theorem 3.1, we provide explicit non-asymptotic high probability convergence rates for the $Q$-learning algorithm presented in Algorithm 1.
Theorem 3.4.
Let $(\widetilde{\gamma}_{t})_{t\in\mathbb{N}}$, and define for all $(x,a)\in\mathcal{X}\times A$ and for all $t\in\mathbb{N}$
(3.9) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}$
Moreover, let $0<\alpha<1$ and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(3.10) $\sum^{\infty}_{t=0}\gamma_{t}(x,a,X_{t})=\infty,\ \sum^{\infty}_{t=0}\gamma_{t}^{2}(x,a,X_{t})<\infty,\ \text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
  • (i)
    For any $\delta\in(0,1)$, and any $Q_{0}\in\mathbb{R}^{|\mathcal{X}|\times|A|}$ it holds with probability [^2] Here we mean w.r.t. the probability measure $\mathbb{P}_{x_{0},\textbf{a}}$. at least $1-\delta$, for all $T\in\mathbb{N}$,
    $\max_{(x,a)\in\mathcal{X}\times A}|Q_{T}(x,a)-Q^{\ast}(x,a)|\;\leq\;$ $\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)\frac{\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)}{{1-\alpha}}$
    $+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\sum_{t=1}^{N_{T}^{*}}\left(\widetilde{\gamma}_{t}\prod_{j=t+1}^{N_{T}^{*}-1}(1-\widetilde{\gamma}_{j})\right)^{2}}$
    $+\frac{C_{r}}{(1-\alpha)^{2}}\cdot\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}^{*}}}$
    where $N_{T}(x,a):=\sum_{s=0}^{T}\mathbf{1}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}$ is the number of visits to $(x,a)\in\mathcal{X}\times A$ up to time $T$, and $N_{T}^{*}=\min_{(x,a)\in\mathcal{X}\times A}N_{T}(x,a)$ denotes the minimal number of visits to all state action pairs until time $T$.
  • (ii)
    With the choice $\widetilde{\gamma}_{t}=\frac{1}{t+1}$ for all $t\in\mathbb{N}$, we have
    $\max_{(x,a)\in\mathcal{X}\times A}|Q_{T}(x,a)-Q^{\ast}(x,a)|\;\leq\;$ $\frac{\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)}{N_{T}^{*}(1-\alpha)}+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{\frac{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)}{N_{T}^{*}}}$
    $+\frac{C_{r}}{(1-\alpha)^{2}}\cdot\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}^{*}}}.$
Note that the error bound in Theorem 3.4 naturally decomposes into three terms. The first two terms correspond to the stochastic approximation error of the QQ–learning recursion. The third term is specific to Algorithm 1 and reflects the statistical error arising from estimating the candidate transition kernels $\mathbb{P}^{(k)}$ via empirical measures. In the special case where the transition kernels are known, this third term vanishes.
4. Approximation of other Ambiguity sets with finite ambiguity sets
In this section we show that using finite ambiguity sets - and hence our proposed $Q$-learning algorithm - provides a versatile framework for ambiguity modeling, as it cannot only be used to model ambiguity w.r.t. a finite pool of candidate models but also to approximate a wide range of other classes of infinitely large ambiguity sets.
4.1. Approximation of infinite ambiguity sets
Consider some possibly infinitely large ambiguity set of probability measures
$\begin{aligned} \mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a)\end{aligned}\subseteq\mathcal{M}_{1}(\mathcal{X})$
for which we want to compute its $Q$-value function $Q^{*}(x,a)=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$. Moreover, we assume that $\mathcal{P}(x,a)$ and the reward function fulfils the following assumption.
Assumption 4.1.
The set-valued map
$\mathcal{X}\times A$ $\rightarrow\mathcal{M}_{1}(\mathcal{X})$
$(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)$
is assumed to be nonempty, compact-valued, and continuous in the weak topology.
The above assumption ensures that minimizers of the corresponding value function exist (compare [38, Theorem 2.7]). Under these assumptions, we can derive the following Lemma 4.2 that allows to approximate the corresponding $Q$-value function by the $Q$-value function of an approximating sequence of finite ambiguity sets as defined in (2.1), justifying in return to apply Algorithm 1 to compute an approximation of the $Q$-value function associated to $\mathcal{P}(x,a)$.
Lemma 4.2.
Assume Assumption 4.1 holds true. If there exists a sequence of finite ambiguity sets $\left(\mathcal{P}^{(n)}(x,a)\right)_{n\in\mathbb{N}}\subset\mathcal{P}(x,a)$ such that for all $(x,a)\in\mathcal{X}\times A$ and all $\mathbb{P}\in\mathcal{P}(x,a)$ there exists a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ for all $n\in\mathbb{N}$ such that [^3] Note that the conditions in (4.1) are equivalent to convergence in the Wasserstein-$1$-distance, compare, e.g. [58, Definition 6.8 (i)], and Section 4.2.
(4.1) $\mathbb{P}^{(n)}\rightarrow\mathbb{P}\text{ weakly, and }\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)\text{ as }n\rightarrow\infty,$
then we also have
$\lim_{n\rightarrow\infty}{Q}^{(n)^{*}}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q^{*}(x,a).$
where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$.
4.2. Approximation of Wasserstein ambiguity sets
Recently, [39] introduced a $Q$-learning algorithm that incorporates model uncertainty by allowing for distributional ambiguity within a Wasserstein-ball around a reference measure. We demonstrate that this method can be approximated by designing a suitable sequence of finite ambiguity sets, and then applying Algorithm 1. Solving the robust $Q$-learning problem for these finite sets then yields an approximation of the optimal $Q$-function corresponding to Wasserstein ambiguity.
For any $q\in\mathbb{N}$ and any probability measures $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{M}_{1}(\mathcal{X})$, the $q$-Wasserstein distance is defined by
$W_{q}(\mathbb{P}_{1},\mathbb{P}_{2}):=\left(\inf_{\pi\in\Pi(\mathbb{P}_{1},\mathbb{P}_{2})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|^{q}\,\mathrm{d}\pi(x,y)\right)^{1/q},$
where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$, and where $\Pi(\mathbb{P}_{1},\mathbb{P}_{2})\subset\mathcal{M}_{1}(\mathcal{X}\times\mathcal{X})$ denotes the set of all couplings of $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$.
Since we work on a finite state space, each measure $\mathbb{P}_{i}$ can be represented as
$\mathbb{P}_{i}=\sum_{x\in\mathcal{X}}a_{i,x}\delta_{x},\qquad\sum_{x\in\mathcal{X}}a_{i,x}=1,\quad a_{i,x}\geq 0,\quad i=1,2,$
where $\delta_{x}$ denotes the Dirac mass at $x\in\mathcal{X}$. In this case, the Wasserstein distance admits the equivalent discrete formulation
$W_{q}(\mathbb{P}_{1},\mathbb{P}_{2})=\left(\min_{\pi_{x,y}\in\widetilde{\Pi}(\mathbb{P}_{1},\mathbb{P}_{2})}\sum_{x,y\in\mathcal{X}}\|x-y\|^{q}\,\pi_{x,y}\right)^{1/q},$
where
$\widetilde{\Pi}(\mathbb{P}_{1},\mathbb{P}_{2}):=\left\{(\pi_{x,y})_{x,y\in\mathcal{X}}\subseteq[0,1]\ \middle|\ \sum_{x^{\prime}\in\mathcal{X}}\pi_{x^{\prime},y}=a_{2,y},\ \sum_{y^{\prime}\in\mathcal{X}}\pi_{x,y^{\prime}}=a_{1,x},\ \forall\,x,y\in\mathcal{X}\right\}.$
Fix $\varepsilon>0$ and $q\in\mathbb{N}$. For each $(x,a)\in\mathcal{X}\times A$, we define the ambiguity set
(4.2) $\mathcal{P}^{(q,\varepsilon)}(x,a):=\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,W_{q}(\mathbb{P},\widehat{\mathbb{P}}(x,a))\leq\varepsilon\right\},$
i.e., the closed $q$-Wasserstein ball of radius $\varepsilon$ centered at some reference distribution $\widehat{\mathbb{P}}(x,a)\in\mathcal{M}_{1}(\mathcal{X})$. We impose the following assumption on the reference kernel.
Assumption 4.3.
The map
$\mathcal{X}\times A$ $\rightarrow\mathcal{M}_{1}(\mathcal{X})$
$(x,a)$ $\mapsto\widehat{\mathbb{P}}(x,a)$
is assumed to be continuous in the Wasserstein-$q$ topology.
Thus, for any state–action pair $(x,a)$, the set $\mathcal{P}^{(q,\varepsilon)}(x,a)$ consists of all probability measures on $\mathcal{X}$ that are within Wasserstein distance $\varepsilon$ of the reference measure $\widehat{\mathbb{P}}(x,a)$.
Lemma 4.4.
Let Assumption 4.3 hold true. Define for $n\in\mathbb{N}$ the set of probability measures on $\mathcal{X}$ with probabilities defined on an equidistant $1/n$-spaced grid via
$G^{(n)}:=\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~\mathbb{P}(y)=\frac{k_{y}}{n}\text{ for some }k_{y}\in\{0,1,\dots,n\}\text{ for all }y\in\mathcal{X}\right\},$
and define for all $n\in\mathbb{N}$ the ambiguity set
(4.3) $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{(q,\varepsilon)}(x,a)\cap G^{(n)}=\left\{\mathbb{P}\in G^{(n)}~\middle|~W_{q}(\mathbb{P},\widehat{\mathbb{P}}(x,a))\leq\varepsilon\right\}.$
  • (i)
    Let $(x,a)\in\mathcal{X}\times A$, and $\mathbb{P}\in\mathcal{P}^{(q,\varepsilon)}(x,a)$. Then, for all $p\in\mathbb{N}$ and for all $n\in\mathbb{N}$ there exists some $\mathbb{P}^{(n)}\in G^{(n)}$ with
    (4.4) $W_{p}(\mathbb{P}^{(n)},\mathbb{P})\leq\frac{\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/p}}{n^{1/p}},$
    for $|\mathcal{X}|$ denoting the number of states of $\mathcal{X}$.
  • (ii)
    We define for $(x,a)\in\mathcal{X}\times A$
    $Q^{(n)^{*}}(x,a)$ $:=\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right],$
    $Q^{*}(x,a)$ $:=\inf_{\mathbb{P}\in\mathcal{P}^{(q,\varepsilon)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
    where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$ defined in (4.3). Then, there exists some $n_{0}$ such that we have for all $n\geq n_{0}$
    $\sup_{(x,a)\in\mathcal{X}\times A}|Q^{*}(x,a)-Q^{(n)^{*}}(x,a)|\leq\frac{C}{n^{1/q}},$
    for
    $C:=\frac{(L_{r}+\alpha L_{V})(\varepsilon+\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q})}{(1-\alpha)}<\infty,$
    and where
    $L_{r}:=\max_{(x,a)\in\mathcal{X}\times\mathcal{A}}\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|r(x,a,y)-r(x,a,y^{\prime})|}{\|y-y^{\prime}\|}<\infty,\qquad L_{V}:=\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|V^{\ast}(y)-V^{\ast}(y^{\prime})|}{\|y-y^{\prime}\|}<\infty.$
Note that the above explicit approximation rates allow in particular to determine the number of finite measures needed to obtain an approximation - given a desired precision.
4.3. Approximation of parametric ambiguity sets
We now turn to parametric uncertainty, i.e., to the case where one specifies a class of potential conditional transition kernels up to the choice of the model parameter $\theta$ which is subject to ambiguity. We define for each state-action pair $(x,a)\in\mathcal{X}\times A$ the ambiguity set
(4.5) $\mathcal{P}^{\Theta}(x,a):=\left\{\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,\theta\in\Theta(x,a)\right\},$
for some state-action dependent parameter set $\Theta(x,a)\subset\mathbb{R}^{d_{\theta}}$ and some parameterized probability measure $\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})$.
Lemma 4.5.
Let $\Theta(x,a)\subset\mathbb{R}^{d}$ be compact for all $(x,a)\in\mathcal{X}\times A$. And assume that for all $(x,a)\in\mathcal{X}\times A$ the map $\Theta(x,a)\ni\theta\mapsto\widehat{\mathbb{P}}(x,a,\theta)$ is continuous in the Wasserstein-1 topology. Then for all $n\in\mathbb{N}$ there exists a set-valued map
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\Theta^{(n)}(x,a)\subset\Theta(x,a)$
which contains finitely many elements, based on which we define for all $n\in\mathbb{N}$ the ambiguity set
(4.6) $\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{\Theta^{(n)}}(x,a)=\left\{\widehat{\mathbb{P}}(x,a,\theta)\in\mathcal{M}_{1}(\mathcal{X})\,\middle|\,\theta\in\Theta^{(n)}(x,a)\right\},$
such that the assumptions of Lemma 4.2 are fulfilled, and we have
$\lim_{n\rightarrow\infty}Q^{(n)^{*}}(x,a)$ $:=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}^{\Theta}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=:Q^{*}(x,a),$
where $V^{(n)}$ denotes the value function as defined in (2.3) associated to the finite ambiguity set $\mathcal{P}^{(n)}(x,a)$ defined in (4.6).
Example 4.6.
Consider the state space $\mathcal{X}=\{1,\dots,N\}$ for some $N\in\mathbb{N}$. Then, for any two functions $\underline{p}:\mathcal{X}\rightarrow\mathbb{R}_{+},\overline{p}:\mathcal{X}\rightarrow\mathbb{R}_{+}$ with $\underline{p}(\cdot)\leq\overline{p}(\cdot)$, the ambiguity set of binomial distributions
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\{\operatorname{Bin}(N,p),p\in[\underline{p}(x),~\overline{p}(x)]\}$
can be approximated in the sense of Lemma 4.5 by
$\mathcal{P}^{(n)}(x,a):=\left\{\operatorname{Bin}(N,p),p\in\{p_{0},\dots,p_{n}\}\text{ with }p_{i}=\underline{p}(x)+\frac{i}{n}\left(\overline{p}(x)-\underline{p}(x)\right),~~i=0,\dots,n\right\}.$
The proof is reported in the appendix. This means, by Lemma 4.5, to approximate the $Q$-value function associated to ambiguity set $\mathcal{P}(x,a)$, we can compute the $Q$-value function associated to ambiguity set $\mathcal{P}^{(n)}(x,a)$ using Algorithm 1 for a sufficiently large $n$.
Under the additional assumption of Lipschitz-continuity of the parameter $\theta$ in the Wasserstein-distance (see Section 4.2), we are able to also derive explicit approximation rates allowing in particular to determine the needed granularity of the covering radius of $\Theta^{(n)}$ for a fixed precision.
Corollary 4.7.
Let $(x,a)\in\mathcal{X}\times\mathcal{A}$, and assume that $\Theta(x,a)\subset\mathbb{R}^{d_{\theta}}$ is compact for all $(x,a)$ and that there exists a constant $L_{\theta}<\infty$ such that the parametrization is uniformly Lipschitz in $W_{1}$, i.e.
(4.7) $W_{1}\!\big(\widehat{\mathbb{P}}(x,a,\theta),\widehat{\mathbb{P}}(x,a,\theta^{\prime})\big)\leq L_{\theta}\|\theta-\theta^{\prime}\|\qquad\text{for all }(x,a)\in\mathcal{X}\times\mathcal{A},\;\theta,\theta^{\prime}\in\Theta(x,a).$
Let $\Theta^{(n)}(x,a)\subset\Theta(x,a)$ be any finite subset and define the finite ambiguity set
$\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{\Theta^{(n)}}(x,a)=\{\widehat{\mathbb{P}}(x,a,\theta)\,:\,\theta\in\Theta^{(n)}(x,a)\}.$
Define the covering radius
(4.8) $\varepsilon_{n}:=\sup_{(x,a)\in\mathcal{X}\times\mathcal{A}}\;\sup_{\theta\in\Theta(x,a)}\;\inf_{\theta^{\prime}\in\Theta^{(n)}(x,a)}\|\theta-\theta^{\prime}\|.$
Then there exists $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$,
(4.9) $\sup_{(x,a)\in\mathcal{X}\times\mathcal{A}}\big|Q^{\ast}(x,a)-Q^{(n)^{*}}(x,a)\big|\;\leq\;\frac{(L_{r}+\alpha L_{V})\,L_{\theta}}{1-\alpha}\;\varepsilon_{n}$
for
$L_{r}:=\max_{(x,a)\in\mathcal{X}\times\mathcal{A}}\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|r(x,a,y)-r(x,a,y^{\prime})|}{\|y-y^{\prime}\|}<\infty,\qquad L_{V}:=\max_{\begin{subarray}{c}y,y^{\prime}\in\mathcal{X}\\ y\neq y^{\prime}\end{subarray}}\frac{|V(y)-V(y^{\prime})|}{\|y-y^{\prime}\|}<\infty.$
5. Continuous state spaces
In many applications of interest the state space is naturally continuous and the assumption of a finite state space imposed in Section 2 is restrictive. In this section we outline how the robust $Q$-learning framework from Section 3 extends to continuous state spaces via function approximation, in the spirit of the classical deep $Q$-learning literature ([6], [35], [36] and [56]).
5.1. Robust Markov decision problems on continuous domains
Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact state space equipped with its Borel $\sigma$-algebra, and let $A$ be a finite action set. We consider a bounded reward function $r:\mathcal{X}\times A\times\mathcal{X}\to\mathbb{R}$ and a discount factor $\alpha\in(0,1)$. As in Section 2, we assume an $(x,a)$-rectangular ambiguity structure; however, in contrast to Section 2 we now allow $\mathcal{X}$ to be continuous.
For each state-action pair $(x,a)\in\mathcal{X}\times\mathcal{A}$ we are given a finite set of candidate transition kernels
(5.1) $\mathcal{P}(x,a)=\Big\{\mathbb{P}^{(1)}(x,a),\dots,\mathbb{P}^{(N)}(x,a)\Big\}\subset\mathcal{M}_{1}(\mathcal{X}).$
We interpret $\mathcal{P}(x,a)$ as the set of admissible conditional distributions of the next state $X_{1}$ given $(X_{0},A_{0})=(x,a)$ and consider the corresponding robust control objective (over an infinite time horizon) under the worst-case model chosen from these kernels.
Given a bounded function $Q:\mathcal{X}\times A\to\mathbb{R}$, define the robust Bellman operator $\mathcal{H}$ by
(5.2) $(\mathcal{H}Q)(x,a):=\min_{k\in\{1,\dots,N\}}\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\Big[r(x,a,X_{1})+\alpha\max_{b\in A}Q(X_{1},b)\Big].$
A robust optimal $Q$-function $Q^{\ast}$ is then characterized (formally) as a fixed point $Q^{\ast}=\mathcal{H}Q^{\ast}$ and the robust value function satisfies $V^{\ast}(x)=\max_{a\in A}Q^{\ast}(x,a)$.
Remark 5.1.
Under standard boundedness assumptions (see [9, Chapter 7]), $\mathcal{H}$ is a contraction w.r.t.  $\|\cdot\|_{\infty}$ with modulus $\alpha$ (the proof is identical to the finite-state case), which motivates iterative methods based on approximating $\mathcal{T}$.
5.2. Function approximation and a robust deep $Q$-learning objective
When $\mathcal{X}$ is continuous, tabular representations of $Q$ are infeasible and we employ a parametric function class
$Q_{\theta}:\mathcal{X}\times A\to\mathbb{R},\qquad\theta\in\mathbb{R}^{p},$
for instance a neural network ([28], [30]). The goal is to fit $Q_{\theta}$ such that it approximately satisfies the robust Bellman equation (5.2).
Similarly to Algorithm 1 (finite-state case), we assume that we can draw samples from each candidate kernel $\mathbb{P}^{(k)}(x,a)$, i.e., we have access to $N$ simulators that, given $(x,a)$, produce independent samples $X_{1}^{(k)}\sim\mathbb{P}^{(k)}(x,a)$.
Let $\theta^{-}$ denote a (possibly delayed) target parameter. For a given $(x,a)$ define the robust TD target
(5.3) $Y_{\theta^{-}}(x,a):=\min_{k\in\{1,\dots,N\}}\mathbb{E}_{X_{1}\sim\mathbb{P}^{(k)}(x,a)}\Big[r(x,a,X_{1})+\alpha\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1},b)\Big].$
In practice, the inner expectation is approximated via Monte-Carlo: for a chosen $M\in\mathbb{N}$,
(5.4) $\widehat{Y}_{\theta^{-}}(x,a):=\min_{k\in\{1,\dots,N\}}\frac{1}{M}\sum_{j=1}^{M}\Big[r(x,a,X_{1,j}^{(k)})+\alpha\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1,j}^{(k)},b)\Big],\qquad X_{1,j}^{(k)}\sim P^{(k)}(x,a).$
Let $\mu$ be a sampling distribution on $\mathcal{X}\times\mathcal{A}$ (e.g. induced by an exploration policy with experience replay ([19])). We propose to minimize the squared Bellman residual:
[…]
Stochastic gradient descent on (5.5) yields a robust analogue of deep $Q$-learning (DQN) for finite ambiguity sets, similar to the approach presented in [30] for Wasserstein ambiguity set.
Remark 5.2 (On the implementation).
  • (a)
    As in standard DQN ([36], [56]), using a slowly updated target parameter $\theta^{-}$ stabilizes training; the robust modification affects only the target computation through the $\min_{k=1,\dots,N}$ in (5.4).
  • (b)
    One may reduce overestimation bias by replacing the target $\max_{b\in\mathcal{A}}Q_{\theta^{-}}(X_{1,j}^{(k)},b)$ with a Double-DQN target, i.e. selecting the maximizing action using the online network and evaluating it using the target network, see [56].
  • (c)
    Relative to non-robust DQN ([36], [56]), the additional cost is linear in $N$, since for each training pair $(x,a)$ one must evaluate NN Monte-Carlo averages and take their minimum.
  • (d)
    If transitions are collected from the real environment under a single unknown kernel, one may still evaluate (5.4) by resimulating next states $X_{1,j}^{(k)}$ using the available simulators for each candidate model $k$.
Section 5.2 provides a practical extension of Algorithm 1 to continuous state spaces by replacing the tabular update with a parametric approximation and by minimizing the robust Bellman residual. Establishing global convergence guarantees for neural-network function approximation is beyond the scope of this paper; nonetheless, the robust modification is structurally minimal and preserves the key feature of our approach: only sampling access to the candidate kernels $\mathbb{P}^{(k)}(x,a)$ is required.
6. Numerical Experiments
In this section we provide two numerical examples comparing Algorithm 1 with other robust and non-robust $Q$-learning algorithms. To apply the numerical method from Algorithm 1, we use for all of the following examples a sequence of learning rates defined by $\widetilde{\gamma}_{t}=\frac{1}{(1+t)^{0.6}}$ for $t\in\mathbb{N}$, a discount factor of $\alpha=0.95$, as well as an $\epsilon$-greedy policy with $\varepsilon$ decreasing over number of iterations, starting with $1$, and decreasing linearly to $0.05$. All the reported experiments are being executed with 10000001~000~000 iterations on a 11th Gen Intel(R) Core(TM) i7-1165G7 @ 2.80GHz (2.80 GHz) - processor. Further details of the implementation can be found under https://github.com/juliansester/FiniteQLearning.
6.1. Coin toss
Inspired by [38, Example 4.1], we consider an agent playing a coin toss game.
At each time step $t\in\mathbb{N}$ the agent observes the result of $10$ coins where an outcome of heads corresponds to $1$, and tails corresponds to 0. The state $X_{t}$ at time $t\in\mathbb{N}$ is then given by the sum of the $10$ coins value, i.e., we have $\mathcal{X}:=\{0,\cdots,10\}$.
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e., $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e., $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\mathdollar$, in contrast if the agent’s bet is wrong she has to pay $1\mathdollar$. The agent also has the possibility not to play. We model this by considering the following reward function:
(6.1) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a1\hskip-3.0pt\mathrm{l}_{\{x<x^{\prime}\}}-a1\hskip-3.0pt\mathrm{l}_{\{x>x^{\prime}\}}-|a|1\hskip-3.0pt\mathrm{l}_{\{x=x^{\prime}\}},$
where the possible actions are given by $A:=\{-1,0,1\}$ with $a=1$ corresponding to betting $X_{t+1}>X_{t}$, $a=0$ to not playing, and a=1a=-1 to betting $X_{t+1}<X_{t}$. In a next step, we define two different ambiguity sets via [^4] We denote by $\operatorname*{Bin}(n,p)$ a binomial distribution with $n$ number of trials and $p$ being the probability of success in a single trial.
(6.2) $\mathcal{P}^{1}(x,a):=$ $\{\operatorname*{Bin}(10,0.5),~\operatorname*{Bin}(10,0.6)\},$
(6.3) $\mathcal{P}^{2}(x,a):=$ $\{\operatorname*{Bin}(10,0.5),~\operatorname*{Bin}(10,0.3)\},$
and we aim at comparing Algorithm 1 with the Algorithm from [38] which does not directly allow to build the same asymmetric sets as in (6.2) and (6.3). To build comparative ambiguity sets with Wasserstein uncertainty we define ambiguity sets such that all probability measures from $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ are contained, respectively. This leads to [^5] Here $W_{1}$ denotes the Wasserstein-$1$ distance, compare for more details, e.g., [58].
(6.4) $\mathcal{P}^{3}(x,a):=$ $\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 1\right\},$
(6.5) $\mathcal{P}^{4}(x,a):=$ $\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 2\right\},$
which are Wasserstein-balls centered at the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [39, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$, but, by construction they contain many additional distributions (e.g. $\operatorname*{Bin}(10,p)$ with intermediate $p$). This highlights the unavoidable over-coverage induced by metric balls.
We train optimal actions for $\mathcal{P}^{1},\mathcal{P}^{2}$ by Algorithm 1, and optimal actions for $\mathcal{P}^{3},\mathcal{P}^{4}$ using the Algorithm from [39]. Upon training, we depict the trained actions in Table 1.
state $X_{t}$ 0 1 2 3 4 5 6 7 8 9 10
$a_{t}^{\mathcal{P}^{1}}(X_{t})$ 11 11 11 11 11 0 0 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{2}}(X_{t})$ 11 11 11 0 0 0 1-1 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{3}}(X_{t})$ 11 11 11 11 0 0 0 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{4}}(X_{t})$ 11 11 11 0 0 0 0 0 1-1 1-1 1-1
$a_{t}^{\rm non-robust}(X_{t})$ 11 11 11 11 11 0 1-1 1-1 1-1 1-1 1-1
Table 1. The trained actions $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$, $a_{t}^{\mathcal{P}^{3}}$, $a_{t}^{\mathcal{P}^{4}}$, and $a_{t}^{\rm non-robust}$ in dependence of the realized state $X_{t}$
We test the profitability of the actions described in Table 1 by playing 100000100~000 rounds of the game according to the trained policies $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$ $a_{t}^{\mathcal{P}^{3}}$ and $a_{t}^{\mathcal{P}^{4}}$. For simulating the 100000100~000 rounds we assume an underlying binomial distribution $\mathbb{P}=\operatorname*{Bin}(10,p_{\rm true})$ with a fixed probability $p_{\rm true}$ for heads which we vary from $0.1$ to $0.9$. We depict the profits of the considered actions in Table 3 and in Figure 1.
Moreover, in Figure 2 we illustrate the convergence of $\max_{(x,a)\in\mathcal{X}\times A}|Q_{t}(x,a)-Q^{*}(x,a)|$ to 0 as the number of itrations $t$ increases. To this end, we study the two ambiguity sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ in the two cases where the transition probabilities are entirely known and are being learned from samples, respectively. We see that in both cases the algorithm converges after a reasonable number of iterations ($\approx 100,000$) up to a small tolerance, while the convergence without known probabilities is slower, accounting for the fact that there is an additional approximation error due to the mismatch between learned probability and true probability, compare also Theorem 3.4.
With the mentioned computational setup, the compute time for $500,000$ iterations was around $3$ minutes for all of the discussed variants.
$p_{\rm true}$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
Robust, $\mathcal{P}^{1}$ 31240-31240 19121-19121 4283-4283 $15161$ 30484{30484} $\mathbf{29857}$ 13212{13212} 10339-10339 29843{-29843}
Robust, $\mathcal{P}^{2}$ $\mathbf{-24424}$ $\mathbf{5231}$ $\mathbf{21051}$ $\mathbf{24613}$ 22998{22998} $11715$ 4984-4984 18896-18896 31185-31185
Robust, $\mathcal{P}^{3}$ 30003-30003 10859-10859 $9843$ $21900$ $25395$ $22587$ $9653$ 10656-10656 30301-30301
Robust, $\mathcal{P}^{4}$ 24528-24528 $4263$ $16771$ $13415$ $9925$ $13366$ $\mathbf{17345}$ $\mathbf{4731}$ $\mathbf{-24452}$
Non-Robust 31101-31101 18251-18251 521-521 $23175$ $\mathbf{35244}$ $23227$ 1387-1387 18421-18421 31024-31024
Table 2. Overall profit of the game described in Section 6.1 in dependence of different trained strategies (described in Table 1) and of the probability distribution $P=\operatorname*{Bin}(10,p_{\rm true})$ of the simulated underlying process. We indicate with bold characters which is the best performing strategy for each choice of $p_{\rm true}$.
Figure 1. Average per step reward of the game described in Section 6.1 in dependence of different trained strategies (described in Table 1) and of the probability distribution $P=\operatorname*{Bin}(10,p_{\rm true})$ of the simulated underlying process, where $p_{\rm true}$ is indicated on the $x$-axis.
Figure 2. The figure shows the convergence of $\max_{(x,a)\in\mathcal{X}\times A}|Q_{t}(x,a)-Q^{*}(x,a)|$ for the two ambiguity sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ for the cases when the transition probabilities are explicitly learned and are being learned from samples, respectively.
For the base case $p_{\rm true}=0.5$ the non-robust strategy is the best performing strategy. The Wasserstein-strategies are outperformed by the robust strategies taking into account either $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$ in all cases for which $p_{\rm true}\leq 0.6$, in particular for the cases $p_{\rm true}=0.3,0.6$ which are by construction contained in the respective ambiguity sets.
The results imply that the choice of the optimal ambiguity sets depends on the scenarios that are deemed possible: If we know that [^6] This would correspond to the case that the dealer in a casino does not tell whether she uses a biased coin (for which we know that $p_{\rm true}=0.6$ for playing the game with an agent or if she uses a fair coin. $p_{\rm true}\in\{0.5,0.6\}$, then one should exactly consider $\mathcal{P}^{1}$, whereas in that situation the set $\mathcal{P}^{3}$ that also covers the case $p_{\rm true}=0.6$ will lead to a suboptimal because too conservative strategy that also covers the case $p_{\rm true}=0.4$ as a possible scenario.
If an applicant however simply has a best guess of a reference measure and considers the cases $p_{\rm true}=0.4$ and $p_{\rm true}=0.6$ both as equally possible, then it is advisable to pursue the Wasserstein-approach outlined in [39]. To summarize the non-robust strategy dominates exactly at the reference model, the finite-robust strategy dominate exactly on the respective ambiguity set, and the Wasserstein robust strategy dominate only when ambiguity is genuinely diffuse.
Example 6.1 thus illustrates how the choice of the ambiguity set fundamentally affects the structure of the optimal policy and the resulting performance. When the ambiguity set is specified as a finite collection of candidate transition kernels, as in $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, the robust $Q$-learning algorithm yields policies that are tailored to protect against a small number of explicitly modelled adverse scenarios. In particular, the resulting strategies exhibit asymmetric betting regions and early abstention in states where the worst-case model renders the expected payoff close to zero.
In contrast, Wasserstein ambiguity sets such as $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$, which are defined as metric balls around a reference distribution, necessarily include a continuum of intermediate distributions that were not explicitly intended to be protected against. This leads to more conservative policies, characterized by wider neutral regions in which the agent refrains from betting. While this conservatism provides robustness against diffuse model misspecification, it may be suboptimal when the modeller has prior knowledge that uncertainty is concentrated on a small number of specific alternatives.
The numerical results confirm this distinction. When the true data-generating process coincides with one of the distributions contained in the finite ambiguity sets (e.g. $p_{\mathrm{true}}\in\{0.5,0.6\}$ for $\mathcal{P}_{1}$), the corresponding finite-robust strategy consistently outperforms both the Wasserstein-robust and the non-robust strategies. Conversely, when uncertainty is genuinely diffuse and symmetric around the reference measure, Wasserstein-based robustness provides a reasonable compromise.
Overall, this example highlights that finite ambiguity sets enable a form of scenario-based robustness, whereas Wasserstein balls implement neighbourhood-based robustness. The former is preferable when the modeller can identify a small set of plausible adverse scenarios, while the latter is appropriate when uncertainty is best described by deviations around a reference model without clear directional structure.
6.2. Portfolio allocation
This example is a variant of [38, Example 4.3] showcasing the use of tailored asymmetrical ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [8], [29], and [59].
We consider the problem of optimally investing in the stock market, and, to this end, encode the return of a stock by $2$ numeric values: either the return is positive ($1$) or negative (1-1). The space of numerically encoded returns is therefore given by $T:=\{-1,1\}$. At each time step $t$ the decision maker can choose to buy the stock or not, or to short sell the stock. The action $a_{t}$ represent this investment decision, where buying the stock is encoded by $1$ and not buying it by 0, whereas short-selling corresponds to 1-1, i.e., we have $A:=\{-1,0,1\}$.
The agent’s investment decision should depend not only on the most recent return but naturally will depend on the current market situation reflecting autocorrelation of financial time series. Therefore the agent relies her investment decision on the last $h=5$ returns and on the previous position. Hence, we consider the state space
$\mathcal{X}:=T^{h}\times A=\{-1,1\}^{h}\times\{-1,0,1\}.$
The reward is given by the function:
$r:\mathcal{X}\times A\times\mathcal{X}$ $\rightarrow\mathbb{R}$
$\bigg((x_{-h+1},\cdots,x_{0},a_{0}),a_{1},(x_{-h+2},\cdots,x_{1},a_{1})\bigg)$ $\mapsto a_{1}\cdot x_{1}-\eta\cdot|a_{1}-a_{0}|,$
i.e., we reward a correct investment decision and penalize an incorrect investment decision. Additionally, the term $\eta\cdot|a_{1}-a_{0}|$ penalizes position changes, accounting for transaction costs. In our experiments we set $\eta=0.1$.
Now, to train the agent, we construct market simulators in the following way. We consider the historical evolution of the (numerically encoded by 1-1 and $1$) daily returns of the underlying stock. This time series is denoted by $(R_{j})_{j=1,...,N}\subset T^{N}$ for some $N\in\mathbb{N}$ labeling the length of the time series. Given some $x=(x_{-h+1},\cdots,x_{0},a_{0})\in\mathcal{X}$ we then sample uniformly the next return from all historical returns sequences with coinciding history [^7] If no such sequence exist we sample uniformly from $T$. $(x_{-h+1},\cdots,x_{0})$. Then, the state transition of the reference measure is partially deterministic (h1h-1 components of subsequent states coincide) and the next state is of the form
$(x_{-h+2},\cdots,x_{0},y,a_{1})\in\mathcal{X}$
where $y$ is the sampled return and $a_{1}$ the executed action.
To apply this construction to real data, we consider daily returns of the stocks of Alphabet, Nvidia, Apple, Microsoft, Amazon, Broadcom, JP Morgan (Tickers: ”AAPL”, ”MSFT”, ”AMZN”, ”GOOGL”,”AVGO”,”JPM”) in the time period from January $2010$ until December $2025$, where we depict the cumulative sum of the signs of the returns in Figure 3.
Figure 3. Cumulative sum of the sign of the returns and the decomposition of the data in training periods and testing periods.
We split the data into a long training period ranging from January $2010$ until December $2021$ and three testing period starting directly thereafter until February $2023$, January $2024$, and November $2025$, respectively. To construct an ambiguity set $\mathcal{P}$ we isolate specific periods of the training data where the market behaved very volatile (Training Period 1), bearish (Training Period 2) or bullish (Training Period 3). These periods vary from stock to stock and are determined by checking the whole training period for that sub-period of $250$ days with 1) the largest standard deviation, 2) the largest increase, and 3) the largest decrease. We denote the probability measures corresponding to these sub-periods by $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3}$, respectively. Moreover, we construct a market simulator (with distribution denoted by $\mathbb{P}_{4}$) by using all training data. Including the accompanying probability measures $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ in an ambiguity set $\mathcal{P}:=\{\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}\}$ then accounts for the fear of the agent to not trade profitably in all considered (difficult) market scenarios, and in normal periods ($\mathbb{P}_{4}$). The choice of these specific training periods is depicted in Figure 3. Using Algorithm 1, we then train different agents using the the market simulators corresponding to the ambiguity set $\mathcal{P}$ and the market simulators for the non-robust state transitions $\mathbb{P}_{i}$, $i=1,2,3,4$, respectively. We emphasize that to use Algorithm 1 we do not need to know $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ explicitly but only need to be able to sample from these measures.
As a benchmark for evaluation, we consider a simple buy-and-hold strategy. The results (mean average reward per trade across all tickers) of the different strategies evaluated in the test period are shown in Table 3, and they show that using a robust strategy indeed enables to avoid losses in all test periods. Note moreover that in the bearish test period 1 the robust strategy in particular outperforms the non-robust agent trained on all data which is likely due to the inclusion of $\mathbb{P}_{3}$ in the ambiguity set $\mathcal{P}$ which is the best performing strategies in this period. Also note that in test periods 2 and 3 it becomes visible that the robust approach due to its construction of a worst-case approach is better suited to avoid losses than to profit from bearish market scenarios.
Test Period 1 Test Period 2 Test Period 3
Days 299 199 500
Negative Returns 1078 635 1573
Positive Returns 1015 758 1927
Average Rewards
Robust ($\mathcal{P}$) 0.0005 0.0168 0.0069
Non-Robust: Training Period 1 ($\mathbb{P}_{1}$) -0.0460 0.0645 0.0150
Non-Robust: Training Period 2 ($\mathbb{P}_{2}$) -0.0436 0.0733 0.0901
Non-Robust: Training Period 3 ($\mathbb{P}_{3}$) 0.0368 -0.0462 -0.0827
Non-Robust: All Data ($\mathbb{P}_{4}$) -0.0397 0.0469 0.0559
Buy and hold -0.0392 0.0989 0.1025
Table 3. A comparison of the average reward per trade in the three test periods depicted in Figure 3 among different strategies and across all tickers. The best-performing strategy is highlighted with bold letters.
7. Proofs and auxiliary results
7.1. Auxiliary results
We start this section by introducing necessary notation. For any function $f:\mathcal{X}\times A\rightarrow\mathbb{R}$, we write
(7.1) $\|f\|_{\infty}:=\sup_{x\in\mathcal{X}}\sup_{a\in A}|f(x,a)|,$
and we define an estimation of the ambiguity set after $t$ steps of Algorithm 1 via the set-valued map
(7.2) $\mathcal{X}\times A\ni(x,a)$ $\twoheadrightarrow\mathcal{P}_{t}(x,a)=\left\{\mathbb{P}_{t}^{(1)}(x,a),\cdots,\mathbb{P}_{t}^{(N)}(x,a)\right\}\subset\mathcal{M}_{1}(\mathcal{X}).$
where $\mathbb{P}_{t}^{(k)}$ is defined in (3.2). Next, we introduce the operators $\mathcal{H}_{t}$ and $\mathcal{H}$ operating on a function $q:\mathcal{X}\times A\rightarrow\mathbb{R}$ being defined by
(7.3) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}_{t}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right]\in\mathbb{R},$
as well as
(7.4) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right]\in\mathbb{R}.$
Moreover, analogue to (2.5), we define the optimal $Q$-value function with respect to the ambiguity set $\mathcal{P}_{t}$ via
(7.5) $\mathcal{X}\times A\ni(x,a)\mapsto Q_{t}^{*}(x,a):=\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right],$
where - by slight abuse of notation - $V$ is defined as in (2.3) with using the ambiguity set $\mathcal{P}_{t}$.
Lemma 7.1.
Let $0<\alpha<1$, and let the ambiguity set $\mathcal{P}_{t}$ be defined in (7.2). Then the following fixed-point equations hold true for the optimal $Q$-value functions defined in (2.5) and (7.5), respectively
$\mathcal{H}_{t}Q_{t}^{*}(x,a)$ $=Q_{t}^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,\text{ for all }t\in\mathbb{N},$
$\mathcal{H}Q^{*}(x,a)$ $=Q^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A.$
Proof.
This follows directly by definition of $Q_{t}^{*}$ and by Proposition 2.1, compare also [39, Proof of Lemma 17]. ∎
Lemma 7.2.
For any maps $q_{i}:\mathcal{X}\times A\rightarrow\mathbb{R},\ i=1,2,$ we have:
$\|\mathcal{H}_{t}q_{1}-\mathcal{H}_{t}q_{2}\|_{\infty}$ $\leq\alpha\|q_{1}-q_{2}\|_{\infty}\text{ for all }t\in\mathbb{N},$
$\|\mathcal{H}q_{1}-\mathcal{H}q_{2}\|_{\infty}$ $\leq\alpha\|q_{1}-q_{2}\|_{\infty}.$
Proof.
The proof follows analogue to [39, Proof of Lemma 18]. ∎
Next, recall the definition
$\mathbb{P}_{t}^{(k)}(x,a)(y)=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}.$
Lemma 7.3.
Let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(7.6) $\sum^{\infty}_{t=0}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}=\infty,\text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Then, we have that
$\mathbb{P}_{t}^{(k)}(x,a)\rightarrow\mathbb{P}^{(k)}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,k\in\{1,\dots,N\}\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
as $t\rightarrow\infty$.
Proof.
Let $k\in\{1,\dots,N\}$ and let $(x,a)\in\mathcal{X}\times A$. First note that the condition in (7.6) ensures that
(7.7) $N_{t}(x,a):=\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}\rightarrow\infty~~\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely as }t\rightarrow\infty.$
Let $\tau_{j}(x,a)$ be the time of the $j$-th visit of $\left((X_{t},a_{t}(X_{t})\right)_{t\in\mathbb{N}}$ to $(x,a)$. Then, we draw $Y_{j}^{(k)}:=X_{\tau_{j}(x,a)+1}^{(k)}\sim\mathbb{P}^{(k)}(x,a)$. By construction $(Y_{j}^{(k)})_{j\in\mathbb{N}}$ is an i.i.d. sequence of random variables with law $\mathbb{P}^{(k)}(x,a)$. For a fixed next state $y\in\mathcal{X}$, define
$\xi_{j}^{(y)}:=1\hskip-3.0pt\mathrm{l}_{\{Y_{j}^{(k)}=y\}},$
and, for $n\in\mathbb{N}$, let $S_{n}^{(y)}:=\sum_{j=1}^{n}\xi_{j}^{(y)}$. By the strong law of large numbers, we get
(7.8) $\frac{S_{n}^{(y)}}{n}\rightarrow\mathbb{E}\left[\xi_{j}^{(y)}\right]=\mathbb{P}^{(k)}(x,a)(y)~~\qquad\mathbb{P}_{x_{0},\textbf{a}}-\text{almost\ surely as }n\rightarrow\infty.$
Now, note that
$\mathbb{P}_{t}^{(k)}(x,a)(y)=\frac{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}),X_{s+1}^{(k)})=(x,a,y)\}}}{\sum_{s=0}^{t}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}}=\frac{S_{N_{t}(x,a)}^{(y)}}{{N_{t}(x,a)}}$
and the assertion follows with (7.7) and (7.8). ∎
Lemma 7.4.
Let $0<\alpha<1$, and let $(x_{0},\textbf{a})\in\mathcal{X}\times\mathcal{A}$ such that
(7.9) $\sum^{\infty}_{t=0}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}=\infty,\text{ for all }(x,a)\in\mathcal{X}\times A,\ \mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
Then, we have that
$\|Q_{t}^{*}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
which implies
${Q}_{t}^{*}(x,a)\rightarrow Q^{*}(x,a)\text{ for all }(x,a)\in\mathcal{X}\times A,\qquad\mathbb{P}_{x_{0},\textbf{a}}\ -\text{almost\ surely}.$
as $t\rightarrow\infty$.
Proof.
We define $C_{r}:=\max_{(x,a,x^{\prime})\in\mathcal{X}\times A\times\mathcal{X}}|r(x,a,x^{\prime})|<\infty$, and then observe that for all $x\in\mathcal{X}$:
$V(x)=\sup_{\mathbf{a}\in\mathcal{A}}\inf_{\mathbb{P}\in\mathfrak{P}_{x,\textbf{a}}}\mathbb{E}_{\mathbb{P}}\left[\sum^{\infty}_{t=0}\alpha^{t}\cdot r(X_{t},a_{t}(X_{t}),X_{t+1})\right]\leq\sum^{\infty}_{t=0}\alpha^{t}\cdot C_{r}=\frac{C_{r}}{1-\alpha}.$
By Proposition 2.1, for all $(x,a)\in\mathcal{X}\times A$ we then obtain
$Q^{*}(x,a)\leq\max_{a\in A}Q^{*}(x,a)=V(x)\leq\frac{C_{r}}{1-\alpha}$
showing that $\|Q^{*}\|_{\infty}<\infty$. Next, for all $(x,a)\in\mathcal{X}\times A$ we compute
(7.10) $|\mathcal{H}_{t}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)|$
$=\left|\min_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]-\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]\right|$
$=\bigg|\min_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\mathbb{P}_{t}^{(k)}(x,a)(y)$
$\hskip 170.71652pt-\min_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\mathbb{P}^{(k)}(x,a)(y)\bigg|$
$\leq\max_{k=1,\dots,N}\bigg|\sum_{y\in\mathcal{X}}\left(r(x,a,y)+\alpha\max_{b\in A}Q^{*}(y,b)\right)\left(\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right)\bigg|$
$\leq\left(C_{r}+\alpha\|Q^{*}\|_{\infty}\right)\max_{k=1,\dots,N}\bigg|\sum_{y\in\mathcal{X}}\left(\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right)\bigg|\rightarrow 0\text{ almost surely, as }t\rightarrow\infty$
where the convergence follows by Lemma 7.3 and since $\|Q^{*}\|_{\infty}<\infty$. Next, we get by using Lemma 7.1 and Lemma 7.2 that
$\|Q_{t}^{*}-Q^{*}\|_{\infty}$ $=\|\mathcal{H}_{t}Q_{t}^{*}-\mathcal{H}Q^{*}\|_{\infty}$
$=\|\mathcal{H}_{t}Q_{t}^{*}-\mathcal{H}_{t}Q^{*}+\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
$\leq\alpha\|Q_{t}^{*}-Q^{*}\|_{\infty}+\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
implying
$\|Q_{t}^{*}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}_{t}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}$
which by (7.10) converges against 0 almost surely as $t\rightarrow\infty$. ∎
The following lemma is a result from stochastic approximation theory compare also [49, Lemma 1], [16] [23, Theorem 1], [51, Lemma 12], and [57, Lemma 3] for further reference.
Lemma 7.5 ([49], Lemma 1).
Let $\mathbb{P}_{0}\in\mathcal{M}_{1}(\Omega)$ be a probability measure on $(\Omega,\mathcal{F})$, and consider a family of stochastic processes $(\gamma_{t}(x,a),F_{t}(x,a),\Delta_{t}(x,a))_{t\in\mathbb{N}_{0}}$, $(x,a)\in\mathcal{X}\times A$, satisfying for all $t\in\mathbb{N}_{0}$
$\Delta_{t+1}(x,a)=\left(1-\gamma_{t}(x,a)\right){\Delta}_{t}(x,a)+\gamma_{t}(x,a)F_{t}(x,a)\qquad\mathbb{P}_{0}\text{-almost surely for all }(x,a)\in\mathcal{X}\times A.$
Let $(\mathcal{G}_{t})_{t\in\mathbb{N}_{0}}\subseteq\mathcal{F}$ be a sequence of increasing $\sigma$-algebras such that for all $(x,a)\in\mathcal{X}\times A$ the random variables $\Delta_{0}(x,a)$ and $\gamma_{0}(x,a)$ are $\mathcal{G}_{0}$-measurable and such that $\Delta_{t}(x,a)$, $\gamma_{t}(x,a)$, and $F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable for all $t\in\mathbb{N}$. Further assume that the following conditions hold.
  • (i)
    $0\leq\gamma_{t}(x,a)\leq 1$, $\sum_{t=0}^{\infty}\gamma_{t}(x,a)=\infty$, $\sum_{t=0}^{\infty}\gamma_{t}^{2}(x,a)<\infty$ $\mathbb{P}_{0}$-almost surely for all $(x,a)\in\mathcal{X}\times A$, $t\in\mathbb{N}_{0}$.
  • (ii)
    There exists $\delta\in(0,1)$ such that $\|\mathbb{E}_{\mathbb{P}_{0}}\left[F_{t}(\cdot,\ \cdot)~\middle|~\mathcal{G}_{t}\right]\|_{\infty}\leq\delta\|\Delta_{t}\|_{\infty}+c_{t}$  $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$, where $c_{t}$ converges to zero  $\mathbb{P}_{0}$-almost surely.
  • (iii)
    There exists $C>0$ such that $\left\|\operatorname{Var}_{\mathbb{P}_{0}}\left(F_{t}(\cdot,\ \cdot)~\middle|~\mathcal{G}_{t}\right)\right\|_{\infty}\leq C(1+\|\Delta_{t}\|_{\infty})^{2}$ $\mathbb{P}_{0}$-almost surely for all $t\in\mathbb{N}_{0}$.
Then, $\lim_{t\rightarrow\infty}\Delta_{t}(x,a)=0\qquad\mathbb{P}_{0}$-almost surely for all $(x,a)\in\mathcal{X}\times A$.
The following result is usually referred to as Popoviciu’s inequality, see [43] or [46].
Lemma 7.6.
Let $Z$ be a random variable on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ satisfying $m\leq Z\leq M$ for some $\infty<m\leq M<\infty$. Then, we have
$\operatorname{Var}_{\mathbb{P}}(Z)\leq\frac{1}{4}\left(M-m\right)^{2}.$
To keep the paper self-contained, we further state the so called Azuma–Hoeffding Inequality, see, e.g., [33] or [34].
Lemma 7.7 ([5]).
Let $(M_{t},\mathcal{F}_{t})_{t\in\mathbb{N}}$ be a martingale with respect to the filtration $(\mathcal{F}_{t})_{t\in\mathbb{N}}$. Assume that the martingale differences are almost surely bounded, i.e., there exist constants $c_{t}>0$ such that
$|M_{t}-M_{t-1}|\leq c_{t}\quad\text{a.s. for all }t\in\mathbb{N}$
Then for any $\varepsilon>0$, $n\in\mathbb{N}$ we have
$\mathbb{P}\!\left(|M_{n}-M_{0}|\geq\varepsilon\right)\leq 2\exp\!\left(-\frac{\varepsilon^{2}}{2\sum_{t=1}^{n}c_{t}^{2}}\right).$
Moreover, we make use of the following concentration inequality for empirical distributions.
Lemma 7.8 ([63], Theorem 2.1).
Let $\mathbb{Q}$ be a probability distribution on $\mathcal{X}$. Let $\widehat{\mathbb{\mathbb{Q}}}_{n}$ denote the empirical distribution of $n$ samples from $\mathbb{Q}$. Then for all $\varepsilon>0$
$\mathbb{P}\left(\|\widehat{\mathbb{\mathbb{Q}}}_{n}-\mathbb{Q}\|_{1}\geq\varepsilon\right)\leq\left(2^{|\mathcal{X}|}-2\right)\exp\left(-\frac{n\varepsilon^{2}}{2}\right).$
7.2. Proofs of the results from Section 2, Section 3, and Section 4
In this section we provide the proofs of Section 2.4 and Section 3.
Proof of Proposition 2.1.
Note that the equality $\max_{a\in A}Q^{*}(x,a)=\mathcal{T}V(x)$ directly follows by definition of the operator $\mathcal{T}$, and it only remains to show $\mathcal{T}V=V$. To this end, we aim at applying [38, Theorem 3.1] with $p=0$ and verify its assumptions, i.e., we verify [38, Assumption 2.2] and [38, Assumption 2.4] by showing the following four properties.
  • (i)
    The set-valued map
    $\mathcal{X}\times A$ $\rightarrow(\mathcal{M}_{1}(\mathcal{X})^{N},\tau_{0})$
    $(x,a)$ $\twoheadrightarrow\mathcal{P}(x,a)=\left\{\mathbb{P}^{(1)}(x,a),\cdots,\mathbb{P}^{(N)}(x,a)\right\}$
    is nonempty, compact-valued and continuous [^8] Continuity of a set-valued map means the map is both upper hemicontinuous and lower hemicontinuous..
  • (ii)
    There exists $C_{P}\geq 1$ such that for all $(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)$ we have
    (7.11) $\int_{X}(1+\|y\|^{p})\mathbb{P}(dy)\leq C_{P}(1+\|x\|^{p})$
  • (iii)
    The reward function $r:\mathcal{X}\times A\times\mathcal{X}\rightarrow\mathbb{R}$ is Lipschitz-continuous in its first two components.
  • (iv)
    We have $0<\alpha<\frac{1}{C_{P}}$, where $C_{P}$ is defined in (ii).
To see that (i) holds, we note that $\mathcal{P}$ is by definition non-empty. It is also compact for all $(x,a)\in\mathcal{X}\times A$ since the set $\mathcal{P}(x,a)$ is finite. To show the upper hemicontinuity, we consider some $(x,a)\in\mathcal{X}\times A$ and we let $(x_{n},a_{n},\mathbb{P}_{n})_{n\in\mathbb{N}}\subseteq\operatorname{Gr}(\mathcal{P}):=\{(x,a,\mathbb{P})~|~(x,a)\in\mathcal{X}\times A,\mathbb{P}\in\mathcal{P}(x,a)\}$ be a sequence such that $(x_{n},a_{n})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$. Since both $\mathcal{X}$ and $A$ are finite, we then have $(x_{n},a_{n})=(x,a)$ for $n$ large enough. Hence, $\mathbb{P}_{n}\in\mathcal{P}(x,a)$ for $n$ large enough. As $\mathcal{P}(x,a)$ is finite we also have that $\mathbb{P}_{n}=\mathbb{P}^{(i)}(x,a)$ for infinitely many indices $n$ for some $i\in\{1,\dots,N\}$. Hence, we can find a subsequence $(\mathbb{P}_{n_{k}})_{k\in\mathbb{N}}$ such that $\mathbb{P}_{n_{k}}\rightarrow\mathbb{P}^{(i)}(x,a)\in\mathcal{P}(x,a)$ weakly as $k\rightarrow\infty$ and the upper hemicontinuity of $\mathcal{P}$ follows by the characterization provided in [2, Theorem 17.20].
To show the lower hemicontinuity let $(x,a)\in\mathcal{X}\times A$, $\mathbb{P}\in\mathcal{P}(x,a)$ and consider a sequence $(x_{n},a_{n})_{n\in\mathbb{N}}\subseteq\mathcal{X}\times A$ with $(x_{n},a_{n})\rightarrow(x,a)$ as $n\rightarrow\infty$. By definition of $\mathcal{P}$, there exists some $i_{\mathbb{P}}\in\{1,\dots,N\}$ such that $\mathbb{P}^{(i_{\mathbb{P}})}(x,a)=\mathbb{P}$. We use this observation to define
$\mathbb{P}_{n}:=\mathbb{P}^{(i_{\mathbb{P}})}(x_{n},a_{n})\in\mathcal{P}(x_{n},a_{n}),~n\in\mathbb{N},$
and we obtain, since $(x_{n},a_{n})=(x,a)$ for $n$ large enough that
$\mathbb{P}_{n}=\mathbb{P}^{(i_{\mathbb{P}})}(x_{n},a_{n})=\mathbb{P}^{(i_{\mathbb{P}})}(x,a)=\mathbb{P}$
for $n$ large enough that and hence $\mathbb{P}_{n}\rightarrow\mathbb{P}$ weakly which shows the lower hemicontinuity of $\mathcal{P}$ according to [2, Theorem 17.21].
To show (ii), note that $p=0$, and define $C_{P}:=1$, then we have
$\int_{\mathcal{X}}(1+\|y\|^{0})\mathbb{P}(dy)=2=C_{P}\cdot(1+\|x\|^{0})$
for all $(x,a)\in\mathcal{X}\times A$. Next, (iii) follows since for all $x_{0},x_{0}^{\prime},x_{1}\in\mathcal{X}$ and $a,a^{\prime}\in A$ we have
$|r(x_{0},a,x_{1})-r(x_{0}^{\prime},a^{\prime},x_{1})|\leq\left(\max_{y_{0},y_{0}^{\prime}\in\mathcal{X},~b,b^{\prime}\in A\atop(y_{0},b)\neq(y_{0}^{\prime},b^{\prime})}\frac{|r(y_{0},b,x_{1})-r(y_{0}^{\prime},b^{\prime},x_{1})|}{\|y_{0}-y_{0}^{\prime}\|+\|b-b^{\prime}\|}\right)\cdot\left(\|x_{0}-x_{0}^{\prime}\|+\|a-a^{\prime}\|\right).$
Finally, (iv) follows since $C_{P}=1$ and as we have $0<\alpha<1$ by assumption. ∎
Proof of Theorem 3.1.
To establish convergence of the $Q$-learning algorithm from Algorithm 1, we apply Lemma 7.5, and to verify the assumptions of Lemma 7.5 we first note that for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ the update rule from Algorithm 1, expressed in (3.3), can be written equivalently as
(7.12) $Q_{t+1}(x,a)=Q_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big);$
where for $(x,a)\in\mathcal{X}\times A$ we recall the definition
(7.13) $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}.$
Next, for each $t\in\mathbb{N}$ and $(x,a)\in\mathcal{X}\times A$ we define
$\Delta_{t}(x,a):$ $=Q_{t}(x,a)-Q_{t}^{*}(x,a),$
$F_{t}(x,a):$ $=\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}.$
Moreover, we recall that as defined in Algorithm 1 we have
$k_{t}^{*}\in\operatorname*{arg\,min}_{k=1,...,N}\mathbb{E}_{\mathbb{P}_{t}^{(k)}(X_{t},a_{t}(X_{t}))}\left[r(X_{t},a_{t}(X_{t}),X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right].$
Once the three assumptions of Lemma 7.5 are verified, it follows that $\Delta_{t}(x,a)\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely as $t\rightarrow\infty$ and hence with Lemma 7.4
$\|Q_{t}-Q^{*}\|_{\infty}\leq\|Q_{t}-Q_{t}^{*}\|_{\infty}+\|Q_{t}^{*}-Q^{*}\|_{\infty}=\|\Delta_{t}\|_{\infty}+\|Q_{t}^{*}-Q^{*}\|_{\infty}\rightarrow\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
and thus
$Q_{t}(x,a)\rightarrow Q^{*}(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$
implying the assertion. Thus, to apply Lemma 7.5, we compute for all $(x,a)\in\mathcal{X}\times A$ that
$\Delta_{t+1}(x,a)=$ $Q_{t}(x,a)-Q_{t}^{*}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big)$
$=$ $\Delta_{t}(x,a)$
$+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)+Q_{t}^{*}(x,a)-Q_{t}^{*}(x,a)\Big)$
$=$ $(1-\gamma_{t}(x,a,X_{t}))\cdot\Delta_{t}(x,a)$
$+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\Big)$
$=$ $(1-\gamma_{t}(x,a,X_{t}))\cdot\Delta_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot F_{t}(x,a),$
where we used in the last step that the indicator $1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}$ appears both in the definition of $F_{t}$ and $\gamma_{t}$ and that the square of the indicator function is the indicator function itself. Consider the filtration $(\mathcal{G}_{t})_{t\in\mathbb{N}}$ with
$\mathcal{G}_{t}:=\sigma(X_{1}^{(1)},X_{2}^{(1)},...,X_{t}^{(N)}),\ t\in\mathbb{N},$
and $\mathcal{G}_{0}:=\{\varnothing,\Omega^{N}\}$ being the trivial sigma-algebra. Then, by definition we have for all $t\in\mathbb{N}$ and for all $(x,a)\in\mathcal{X}\times A$ that the random variables $\Delta_{t}(x,a),\ \gamma_{t}(x,a),\ F_{t-1}(x,a)$ are $\mathcal{G}_{t}$-measurable.
We now define the random operator
(7.14) $\mathcal{X}\times A\ni(x,a)\mapsto\widetilde{\mathcal{H}}_{t}q(x,a):=\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right],$
where $k_{t}^{*}$ is the index selected by the empirical minimization in Algorithm 1; expectations are taken under the corresponding true kernel $\mathbb{P}^{(k_{t}^{*})}$ to quantify the kernel estimation error between $\mathbb{P}_{t}^{(k_{t}^{*})}$ and $\mathbb{P}^{(k_{t}^{*})}$. Further, we define $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and then, analogous to the proof of Lemma 7.4, we obtain $\|Q_{t}\|_{\infty}<\frac{C_{r}}{1-\alpha}<\infty$. Next, let $(x,a)\in\mathcal{X}\times A,$ and $t\in\mathbb{N}$, and note that
$\left|\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-{\mathcal{H}}_{t}Q_{t}(x,a)\right|$
$=\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]-\mathbb{E}_{\mathbb{P}_{t}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right|$
$\leq\max_{k=1,\dots,N}\left|\mathbb{E}_{\mathbb{P}^{(k)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]-\mathbb{E}_{\mathbb{P}_{t}^{(k)}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]\right|$
$\leq\left(C_{r}+\alpha\|Q_{t}\|_{\infty}\right)\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\sum_{y\in\mathcal{X}}\left|\mathbb{P}_{t}^{(k)}(x,a)(y)-\mathbb{P}^{(k)}(x,a)(y)\right|=:c_{t}$
with $c_{t}\rightarrow 0$ $\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$ by Lemma 7.3 and since $\|Q_{t}\|_{\infty}<\infty$. Then, we have by the definition of $\mathbb{P}_{x_{0},\textbf{a}}$ in (3.4) $\mathbb{P}_{x_{0},\textbf{a}}$-almost surely that
$|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)~\middle|~\mathcal{G}_{t}\right]\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathbb{E}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right]\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)\right|$
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left|\mathcal{H}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)+\widetilde{\mathcal{H}}_{t}Q_{t}(x,a)-{\mathcal{H}}_{t}Q_{t}(x,a)\right|$
$\leq 1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left(\left|\mathcal{H}_{t}Q_{t}(x,a)-Q_{t}^{*}(x,a)\right|+c_{t}\right)$
Consequently, applying Lemma 7.1 we have
(7.15) $|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(x,a)|\mathcal{G}_{t}]|\leq 1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\left(|\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}^{*}(x,a)|+c_{t}\right).$
Now using (7.15) and applying Lemma 7.2 we can conclude
$\|\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}[F_{t}(\cdot,\cdot)|\mathcal{G}_{t}]\|_{\infty}\leq\|\mathcal{H}_{t}Q_{t}-\mathcal{H}_{t}Q_{t}^{*}\|_{\infty}+c_{t}\leq\alpha\|Q_{t}-Q_{t}^{*}\|_{\infty}+c_{t}=\alpha\|\Delta_{t}\|_{\infty}+c_{t}$
showing Lemma 7.5 (ii).
Next, to show that the assumption from Lemma 7.5 (iii) is fulfilled, we observe that for all $(x,a)\in\mathcal{X}\times A$ and $t\in\mathbb{N}$ we have
[…]
$=1\hskip-3.0pt\mathrm{l}_{\{(x,a)=(X_{t},~a_{t}(X_{t}))\}}\cdot\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)$
$\leq\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}^{*}(x,a)\right)\right)$
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right)$
$=\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})\right).$
We recall $C_{r}:=\max_{y_{0},y_{1}\in\mathcal{X},\ b\in A}|r(y_{0},b,y_{1})|<\infty$ and compute as an upper bound for the integrand from (7.16) that
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\leq C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\max_{b^{\prime}\in A}Q_{t}^{*}(X_{t+1},b^{\prime})$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}(\max_{b^{\prime}\in A}Q_{t}^{*}(y,b^{\prime})-\max_{b\in A}Q_{t}(y,b))$
$\leq C_{r}+\alpha\max_{y\in\mathcal{X}}\max_{b\in A}|Q_{t}^{*}(y,b)-Q_{t}(y,b)|$
$\leq C_{r}+\alpha\|\Delta_{t}\|_{\infty}=:M\in\mathbb{R}.$
On the other hand we also may compute a lower bound as follows
$-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})$
$\geq-C_{r}+\alpha(\min_{y^{\prime}\in\mathcal{X}}\min_{b\in A}(Q_{t}^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)))$
$\geq-C_{r}+\alpha(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y^{\prime},b)-Q_{t}(X_{t+1},b)|)$
$\geq-C_{r}+\alpha\min_{y\in\mathcal{X}}(-\max_{y^{\prime}\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y^{\prime},b)-Q_{t}(y,b)|)$
$\geq-C_{r}-\alpha(\max_{y^{\prime},y\in\mathcal{X},\ b\in A}|Q_{t}(y,b)-Q_{t}^{*}(y,b)|+|Q_{t}^{*}(y,b)-Q_{t}^{*}(y^{\prime},b)|)$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\max_{(y^{\prime},y)\in\mathcal{X},\ b\in A}|Q_{t}^{*}(y,b)-Q_{t}^{*}(y^{\prime},b)|$
$\geq-C_{r}-\alpha\|\Delta_{t}\|_{\infty}-\alpha\frac{2C_{r}}{1-\alpha}=:m\in\mathbb{R}.$
We apply Popoviciu’s inequality on variances from Lemma 7.6 with the bounds $m$ and $M$ computed just above, to (7.16) and obtain
$\operatorname{Var}_{\mathbb{P}_{x_{0},\textbf{a}}}$ $(F_{t}(x,a)|\mathcal{G}_{t})$
$\leq$ $\operatorname{Var}_{\mathbb{P}^{(k_{t}^{*})}(x,a)}\left(-r(x,a,X_{t+1})-\alpha\max_{b\in A}Q_{t}(X_{t+1},b)+\alpha\min_{y^{\prime}\in\mathcal{X}}\max_{b^{\prime}\in A}Q_{t}^{*}(y^{\prime},b^{\prime})\right)$
$\leq$ $\frac{1}{4}(M-m)^{2}$
$\leq$ $\frac{1}{4}\Big(C_{r}+\alpha\|\Delta_{t}\|_{\infty}+C_{r}+\alpha\|\Delta_{t}\|_{\infty}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}$
$=$ $\frac{1}{4}\Big(C_{r}+C_{r}+2\alpha\|\Delta_{t}\|_{\infty}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}$
$\leq$ $\frac{1}{2}\left(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\left(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}\right)$
$\leq$ $\Bigg(\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}+4\alpha^{2}\|\Delta_{t}\|_{\infty}^{2}+4\alpha^{2}+\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}\|\Delta_{t}\|_{\infty}^{2}\Bigg)$
$\leq$ $\left(4\alpha^{2}+\Big(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha}\Big)^{2}\right)(1+\|\Delta_{t}\|_{\infty}^{2})\leq C\cdot(1+\|\Delta_{t}\|_{\infty})^{2},$
with
(7.17) $C:=\left(4\alpha^{2}+(2C_{r}+\alpha\frac{2C_{r}}{1-\alpha})^{2}\right).$
This shows that also Lemma 7.5 (iii) is fulfilled, and hence, with an application of Lemma 7.5 we obtain $Q_{t}(x,a)\rightarrow Q(x,a)\quad\mathbb{P}_{x_{0},\textbf{a}}\text{-almost\ surely as }t\rightarrow\infty$ for all $(x,a)\in\mathcal{X}\times A$. ∎
Proof of Theorem 3.4.
We start by showing the claim of (i).
To this end, we first decompose the difference between the learned $Q$-value function and the optimal $Q$-value function by applying the triangle inequality
(7.18) $\max_{x\in\mathcal{X},a\in A}|Q_{T}(x,a)-Q^{*}(x,a)|=\|Q_{T}-Q^{*}\|_{\infty}\leq\|Q_{T}-Q^{*}_{T}\|_{\infty}+\|Q^{*}_{T}-Q^{*}\|_{\infty}.$
This means to find an upper bound for $\max_{x\in\mathcal{X},a\in A}|Q_{T}(x,a)-Q^{*}(x,a)|$, it suffices to bound the two terms on the right-hand side of (7.18). Note that by Lemma 7.4 we have
$\|Q^{*}_{T}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}_{T}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}.$
Next, we first apply (7.10), and second we observe $\|Q^{*}\|_{\infty}\leq\|V^{*}\|_{\infty}\leq\sum_{t=0}^{\infty}\alpha^{t}C_{r}=\frac{C_{r}}{1-\alpha}$, to get
$\frac{\|\mathcal{H}_{T}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}$ $\leq\frac{C_{r}+\alpha\|Q^{*}\|_{\infty}}{1-\alpha}\cdot\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}$
$\leq\frac{C_{r}+\alpha\frac{C_{r}}{1-\alpha}}{1-\alpha}\cdot\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}$
$=\frac{C_{r}}{(1-\alpha)^{2}}\cdot\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}.$
Then, note that we can bound the difference between the empirical distribution $\mathbb{P}_{t}^{(k)}$ and $\mathbb{P}^{(k)}$ by Lemma 7.8. We obtain for all $T\in\mathbb{N}$ that
$\mathbb{P}_{x_{0},\textbf{a}}\!\left(\bigcap_{(x,a)\in\mathcal{X}\times A,\atop k\in\{1,\dots,N\}}\left\{\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}<\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}(x,a)}}\right\}\!\right)$
$=1-\mathbb{P}_{x_{0},\textbf{a}}\!\left(\bigcup_{(x,a)\in\mathcal{X}\times A,\atop k\in\{1,\dots,N\}}\left\{\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}\geq\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}(x,a)}}\right\}\!\right)$
$\geq 1-\sum_{(x,a)\in\mathcal{X}\times A,\atop k\in\{1,\dots,N\}}\mathbb{P}_{x_{0},\textbf{a}}\!\left(\left\|\mathbb{P}_{T}^{(k)}(x,a)-\mathbb{P}^{(k)}(x,a)\right\|_{1}\geq\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}(x,a)}}\!\right)$
$\geq 1-N|\mathcal{X}||A|\cdot\frac{\delta}{N|\mathcal{X}||A|}=1-\delta$
for $N_{T}(x,a):=\sum_{s=0}^{T}1\hskip-3.0pt\mathrm{l}_{\{(X_{s},a_{s}(X_{s}))=(x,a)\}}$ denoting the number of visits until time $T$ of the state action pair $(x,a)\in\mathcal{X}\times A$. Thus, with probability at least $1-\delta$ we have
(7.19) $\|Q^{*}_{T}-Q^{*}\|_{\infty}$ $\leq\frac{C_{r}}{(1-\alpha)^{2}}\cdot\max_{(x,a)\in\mathcal{X}\times A}\max_{k=1,\dots,N}\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{N_{T}(x,a)}}$
$=\frac{C_{r}}{(1-\alpha)^{2}}\cdot\sqrt{\frac{2\log\left((2^{|\mathcal{X}|}-2)\cdot\left(\frac{N|\mathcal{X}||A|}{\delta}\right)\right)}{\min_{(x,a)\in\mathcal{X}\times A}N_{T}(x,a)}}.$
Next, we turn to the term $\|Q_{T}-Q_{T}^{*}\|_{\infty}$ appearing on the right-hand side of (7.18). Employing the notations $\Delta_{t}:=Q_{t}-Q_{t}^{*}$ and $\gamma_{t}(x,a,X_{t}):=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\widetilde{\gamma}_{N_{t}(x,a)}$ we have by the updating rule of $Q_{t}$ that for all $(x,a)\in\mathcal{X}\times A$
(7.20) $\Delta_{t+1}(x,a)$ $=\Delta_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-Q_{t}(x,a)\Big)$
$=\Delta_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-\Delta_{t}(x,a)-Q_{t}^{*}(x,a)\Big)$
$=\left(1-\gamma_{t}(x,a,X_{t})\right)\Delta_{t}(x,a)$
$\hskip 28.45274pt+\gamma_{t}(x,a,X_{t})\cdot\Big(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-\mathcal{H}_{t}Q_{t}^{*}(x,a)\Big)$
where we used Lemma 7.1 in the last step. We next define
$\varepsilon_{t+1}:=1\hskip-3.0pt\mathrm{l}_{\left\{(x,a)=\left(X_{t},~a_{t}(X_{t})\right)\right\}}\cdot\left(r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)-\mathcal{H}_{t}Q_{t}(x,a)\right),$
and obtain with (7.20) that
(7.21) $\Delta_{t+1}(x,a)$ $=\left(1-\gamma_{t}(x,a,X_{t})\right)\Delta_{t}(x,a)+\gamma_{t}(x,a,X_{t})\cdot\Big(\varepsilon_{t+1}+\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}^{*}(x,a)\Big).$
For $(x,a)\in\mathcal{X}\times A$ let $\tau_{1}(x,a)<\tau_{2}(x,a)<\dots<\tau_{N_{T}(x,a)}(x,a)\leq T$ denote the visit times of $(x,a)$ up to $T$. Then for $j\in\{1,\dots,N_{T}(x,a)\}$ we have at time $t=\tau_{j}(x,a)$ that $\gamma_{t}(x,a,X_{t})=\widetilde{\gamma}_{N_{t}(x,a)}=\widetilde{\gamma}_{j}$ and hence
(7.22) $\Delta_{t+1}(x,a)$ $=\left(1-\widetilde{\gamma}_{j}\right)\Delta_{t}(x,a)+\widetilde{\gamma}_{j}\cdot\Big(\varepsilon_{t+1}+\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}^{*}(x,a)\Big)$
$=\left(1-\widetilde{\gamma}_{j}\right)\Delta_{t}(x,a)+\widetilde{\gamma}_{j}\cdot\xi_{t+1}$
for
$\xi_{t+1}:=\varepsilon_{t+1}+\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}^{*}(x,a).$
Then, unrolling (7.22) along all $t$, is the same as unrolling (7.22) along the subsequence $(\tau_{j}(x,a))_{j}$ as for $t\not\in\left\{\tau_{i}(x,a),i=1,\dots,N_{T}(x,a)\right\}$ it holds $\Delta_{t+1}(x,a)=\Delta_{t}(x,a)$. This gives
$\Delta_{\tau_{N_{T}(x,a)}}(x,a)=\left(\prod_{i=1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{i}\right)\right)\Delta_{0}(x,a)+\sum_{i=1}^{N_{T}(x,a)-1}\left(\widetilde{\gamma}_{i}\xi_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\right).$
Taking absolutes values and applying Lemma 7.2 leads to
$|\Delta_{\tau_{N_{T}(x,a)}}(x,a)|\leq$ $\left(\prod_{i=1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{i}\right)\right)|\Delta_{0}(x,a)|$
$+\bigg|\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\varepsilon_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg|+\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\alpha\max_{t\leq T}\|\Delta_{t}\|_{\infty}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right).$
Now take the maximum over $(x,a)\in\mathcal{X}\times A$ and set $B_{T}:=\max_{t\leq T}\|\Delta_{t}\|_{\infty}$ to obtain
(7.23) $B_{T}\leq$ $\max_{(x,a)\in\mathcal{X}\times A}\bigg\{\left(\prod_{i=1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{i}\right)\right)\|\Delta_{0}\|_{\infty}$
$\hskip 42.67912pt+\bigg|\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\varepsilon_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg|+\alpha B_{T}\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg\}$
$\leq$ $\max_{(x,a)\in\mathcal{X}\times A}\bigg\{\left(\prod_{i=1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{i}\right)\right)\|\Delta_{0}\|_{\infty}\bigg\}$
$\hskip 42.67912pt+\max_{(x,a)\in\mathcal{X}\times A}\bigg\{\bigg|\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\varepsilon_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg|\bigg\}+\alpha B_{T},$
where we used $\sum_{i=0}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)=1-\prod_{j=1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\leq 1$. Hence, we get using (7.23) that
(7.24) $B_{T}$ $\leq\frac{\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)\|\Delta_{0}\|_{\infty}+\max_{(x,a)\in\mathcal{X}\times A}\bigg|\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\varepsilon_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg|}{1-\alpha},$
where $N_{T}^{*}=\min_{(x,a)\in\mathcal{X}\times A}N_{T}(x,a)$, since $\mathbb{N}\ni n\mapsto\prod_{i=1}^{n-1}\left(1-\widetilde{\gamma}_{i}\right)$ is decreasing in $n$. Moreover, we have
$\|\Delta_{0}\|_{\infty}\leq\|Q_{0}\|_{\infty}+\|Q^{*}_{0}\|_{\infty}\leq\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha},$
and thus (7.24) writes as
(7.25) $B_{T}$ $\leq\frac{\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)+\max_{(x,a)\in\mathcal{X}\times A}\bigg|\sum_{i=1}^{N_{T}(x,a)-1}\widetilde{\gamma}_{i}\varepsilon_{\tau_{i}(x,a)+1}\prod_{j=i+1}^{N_{T}(x,a)-1}\left(1-\widetilde{\gamma}_{j}\right)\bigg|}{1-\alpha}.$
Next note that for each fixed $(x,a)\in\mathcal{X}\times A$ and for any sequence $(w_{i})_{i\in\mathbb{N}}\subseteq[0,1]$, the process $(M_{t})_{t\in\mathbb{N}}$ defined by
$M_{t}:=\sum_{s=0}^{t-1}w_{s}\varepsilon_{s+1}.$
forms a martingale w.r.t. the filtration generated by $(X_{t})_{t\in\mathbb{N}}$, denoted by $(\mathcal{F}_{t})_{t\in\mathbb{N}}$ whenever $w_{t}\in\mathcal{F}_{t}$ for all $t\in\mathbb{N}$. Indeed, we have $M_{t+1}=M_{t}$ if $(x,a)\neq(X_{t},a_{t}(X_{t}))$, and in the case $(x,a)=(X_{t},a_{t}(X_{t}))$ we compute
(7.26) $\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[M_{t+1}~\middle|~\mathcal{F}_{t}\right]$ $=M_{t}+w_{t}\left(\mathbb{E}_{\mathbb{P}_{x_{0},\textbf{a}}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\middle|~\mathcal{F}_{t}\right]-\mathcal{H}_{t}Q_{t}(x,a)\right)$
$=M_{t}+w_{t}\left(\inf_{\mathbb{P}\in\mathcal{P}_{t}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{t+1})+\alpha\max_{b\in A}Q_{t}(X_{t+1},b)\right]-\mathcal{H}_{t}Q_{t}(x,a)\right)$
$=M_{t}+w_{t}\left(\mathcal{H}_{t}Q_{t}(x,a)-\mathcal{H}_{t}Q_{t}(x,a)\right)=M_{t}.$
Additionally the increments of $M_{t}$ are bounded:
(7.27) $\left|M_{t+1}-M_{t}\right|=\left|w_{t}{\varepsilon}_{t+1}\right|$ $\leq w_{t}\left(C_{r}+\alpha\max_{x\in\mathcal{X},b\in A}|Q_{t}(x,b)|+\max_{x\in\mathcal{X},a\in A}|\mathcal{H}_{t}Q_{t}(x,a)|\right)$
$\leq w_{t}\left(C_{r}+\alpha\frac{C_{r}}{1-\alpha}+C_{r}+\alpha\frac{C_{r}}{1-\alpha}\right)=w_{t}\frac{2C_{r}}{1-\alpha}.$
As $(M_{t})_{t\in\mathbb{N}}$ is a martingale with bounded increments, by Lemma 7.7 we get for any $u>0$ that
(7.28) $\mathbb{P}_{x_{0},\textbf{a}}\Big(\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|\geq u\Big)\leq 2\exp\Bigg(-\frac{u^{2}}{2(2C_{r}/(1-\alpha))^{2}\sum_{t=1}^{N_{T}(x,a)}w_{t}^{2}}\Bigg).$
Choose
$u(x,a):=\frac{2C_{r}}{1-\alpha}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\sum_{t=1}^{N_{T}(x,a)}w_{t}^{2}}.$
Then, for each $(x,a)\in\mathcal{X}\times A$, inequality (7.28) implies that
(7.29) $\mathbb{P}_{x_{0},\textbf{a}}\Big(\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|\geq u(x,a)\Big)\leq\frac{\delta}{|\mathcal{X}||A|}.$
We now specify
$w_{t}=\widetilde{\gamma}_{t}\prod_{j=t+1}^{N_{T}(x,a)-1}(1-\widetilde{\gamma}_{j})\in[0,1],\qquad w_{t}^{(n)}=\widetilde{\gamma}_{t}\prod_{j=t+1}^{n-1}(1-\widetilde{\gamma}_{j})\in[0,1].$
Since the weights $w_{t}$ depend on the visit count $N_{T}(x,a)$ which is not contained in $\mathcal{F}_{t}$ for $t<T$, we condition on the event $\{N_{T}(x,a)=n\}$. On this event we have $w_{t}=w_{t}^{(n)}$, and the weights are deterministic, and the noise sequence remains a martingale difference. Applying Azuma–Hoeffding conditionally and unconditioning yields the stated bound.
(7.30) $\mathbb{P}_{x_{0},\textbf{a}}\Big(\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|\geq u(x,a)\Big)$
$=\sum_{n=0}^{T}\mathbb{P}_{x_{0},\textbf{a}}\Big(\Big|\sum_{j=1}^{n}w_{j}^{(n)}\varepsilon_{\tau_{j}(x,a)+1}\geq u(x,a)\Big|N_{T}(x,a)=n\Big)\mathbb{P}_{x_{0},\textbf{a}}(N_{T}(x,a)=n)$
$\leq\sum_{n=0}^{T}\frac{\delta}{|\mathcal{X}||A|}\mathbb{P}_{x_{0},\textbf{a}}(N_{T}(x,a)=n)=\frac{\delta}{|\mathcal{X}||A|}.$
We now use (7.30) to obtain
$\mathbb{P}_{x_{0},\textbf{a}}\!\left(\bigcap_{(x,a)\in\mathcal{X}\times A}\left\{\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|<u(x,a)\right\}\!\right)$
$=1-\mathbb{P}_{x_{0},\textbf{a}}\!\left(\bigcup_{(x,a)\in\mathcal{X}\times A}\left\{\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|\geq u(x,a)\right\}\!\right)$
$\geq 1-\sum_{(x,a)\in\mathcal{X}\times A}\mathbb{P}_{x_{0},\textbf{a}}\!\left(\left\{\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|\geq u(x,a)\right\}\!\right)$
$\geq 1-|\mathcal{X}||A|\cdot\frac{\delta}{|\mathcal{X}||A|}=1-\delta.$
This implies that with probability at least $1-\delta$,
(7.31) $\max_{(x,a)\in\mathcal{X}\times A}\Big|\sum_{j=1}^{N_{T}(x,a)}w_{j}\varepsilon_{\tau_{j}(x,a)+1}\Big|$ $\leq\max_{(x,a)\in\mathcal{X}\times A}\frac{2C_{r}}{1-\alpha}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\sum_{t=1}^{N_{T}(x,a)}w_{t}^{2}}$
$=\frac{2C_{r}}{1-\alpha}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\max_{(x,a)\in\mathcal{X}\times A}\sum_{t=1}^{N_{T}(x,a)}w_{t}^{2}}.$
Now combining (7.25) and (7.31) gives with probability at least $1-\delta$ that
(7.32) $\|Q_{T}-Q_{T}^{*}\|_{\infty}$
$=\|\Delta_{T}\|_{\infty}\leq\frac{\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)}{{1-\alpha}}+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\max_{(x,a)\in\mathcal{X}\times A}\sum_{t=1}^{N_{T}(x,a)}w_{t}^{2}}$
$=\frac{\left(\|Q_{0}\|_{\infty}+\frac{C_{r}}{1-\alpha}\right)\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)}{{1-\alpha}}+\frac{2C_{r}}{(1-\alpha)^{2}}\sqrt{2\log\Big(\frac{2|\mathcal{X}||A|}{\delta}\Big)\sum_{t=1}^{N_{T}^{*}}\left(\widetilde{\gamma}_{t}\prod_{j=t+1}^{N_{T}^{*}-1}(1-\widetilde{\gamma}_{j})\right)^{2}}.$
Eventually, we show the assertion of (ii), where with the choice $\widetilde{\gamma}_{t}=\frac{1}{t+1}$ we have
$\prod_{i=1}^{N_{T}^{*}-1}\left(1-\widetilde{\gamma}_{i}\right)=\prod_{i=1}^{N_{T}^{*}-1}\frac{i}{i+1}=\frac{1}{N_{T}^{*}}$
as well as
$\sum_{t=1}^{N_{T}^{*}}w_{t}^{2}=\sum_{t=1}^{N_{T}^{*}}\left(\frac{1}{t+1}\prod_{j=t+1}^{N_{T}^{*}-1}\frac{j}{j+1}\right)^{2}=\frac{N_{T}^{*}-1}{(N_{T}^{*})^{2}}+\frac{1}{(N_{T}^{*}+1)^{2}}\leq\frac{1}{N_{T}^{*}}$
and thus (7.32) simplifies to
[…]
Proof of Lemma 4.2.
Let $(x,a)\in\mathcal{X}\times A$, then, since Assumption 4.1 holds, by [38, Theorem 2.7 (i)], there exists some minimizing measure $\mathbb{P}^{*}\in\mathcal{P}(x,a)$ such that
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]=\mathbb{E}_{\mathbb{P}^{*}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
Take some sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ for all $n\in\mathbb{N}$ such that $\mathbb{P}^{(n)}\rightarrow\mathbb{P}^{*}$ weakly as $n\rightarrow\infty$ and such that $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. Then, since $\mathcal{X}$ is finite, it follows directly that
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$ $=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
$\geq\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
But note that, since $\mathcal{P}^{(n)}(x,a)\subset\mathcal{P}(x,a)$, we also have
$\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]\leq\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
showing that
(7.34) $\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$ $=\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
$=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right].$
Let $V_{\varepsilon_{n}}(x)$ denote the value function defined as in (2.3) associated to the ambiguity set $\mathcal{P}_{\varepsilon_{n}}(x,a):=\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~|~W_{1}(\mathbb{P},\mathbb{P}^{*})\leq\varepsilon_{n}\}$ where $\varepsilon_{n}=W_{1}(\mathbb{P}^{(n)},\mathbb{P}^{*})$, for $W_{1}$ denoting the $1$-Wasserstein-distance defined by
$W_{1}(\mathbb{P}^{(n)},\mathbb{P}^{*}):=\inf_{\pi\in\Pi(\mathbb{P}^{(n)},\mathbb{P}^{*})}\int_{\mathcal{X}\times\mathcal{X}}\|x-y\|{\mathop{}\!\mathrm{d}}\pi(x,y),$
where $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$, and where $\Pi(\mathbb{P}^{(n)},\mathbb{P}^{*})$ denotes the set of joint distributions of two probability measures $\mathbb{P}^{(n)}$ and $\mathbb{P}^{*}$, compare also, e.g. [58]. Then, [40, Theorem 3.1] implies for all $x\in\mathcal{X}$ that
(7.35) $|V^{*}(x)-V^{(n)}(x)|\leq|V^{*}(x)-V_{\varepsilon_{n}}(x)|+|V_{\varepsilon_{n}}(x)-V^{(n)}(x)|\leq C\varepsilon_{n}$
for some constant $C$ independent of $x\in\mathcal{X}$, since both $\mathbb{P}^{*}$ and $\mathbb{P}^{(n)}$ are by definition inside a Wasserstein-ball centered at $\mathbb{P}^{*}$ with radius $\varepsilon_{n}$. Hence, we get together with (7.34)
$\lim_{n\rightarrow\infty}\inf_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V^{(n)}(X_{1})\right]$
$=\lim_{n\rightarrow\infty}\mathbb{E}_{\mathbb{P}^{(n)}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]+\mathbb{E}_{\mathbb{P}^{(n)}}\left[\alpha\left(V^{(n)}(X_{1})-V(X_{1})\right)\right]$
$=\inf_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha V(X_{1})\right]$
by (7.35) since $\varepsilon_{n}=W_{1}(\mathbb{P}^{(n)},\mathbb{P}^{*})\rightarrow 0$ as $n\rightarrow\infty$ which follows from [58, Definition 6.8 (i)] and by the assumption that $\mathbb{P}^{(n)}\rightarrow\mathbb{P}$ weakly as well as $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. ∎
Proof of Lemma 4.4.
Let $\varepsilon>0,q\in\mathbb{N}$, $(x,a)\in\mathcal{X}\times A$, and let $\mathbb{P}\in\mathcal{P}^{(q,\varepsilon)}(x,a)$.
  • (i)
    First, we show that for all $p\in\mathbb{N}$ and for all $n\in\mathbb{N}$ there exists some $\mathbb{P}^{(n)}\in G^{(n)}$ with
    (7.36) $W_{p}(\mathbb{P}^{(n)},\mathbb{P})\leq\frac{C}{n^{1/p}},$
    where $C:=\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/p}$ depends on $p$ but not on $n$.
    To this end, let $n\in\mathbb{N}$ and define for all $y\in\mathcal{X}$ the quantity $k_{y}^{\prime}:=\lfloor n\cdot\mathbb{P}(y)\rfloor\in\{0,1,\dots,n\}$, and select an arbitrary subset $\mathcal{X}^{\prime}\subset\mathcal{X}$ with $|\mathcal{X}^{\prime}|=n-\sum_{y\in\mathcal{X}}k_{y}^{\prime}$. Based on this, we set for all $y\in\mathcal{X}$
[…]
    Then, by definition,
    $\sum_{y\in\mathcal{X}}k_{y}=\sum_{y\in\mathcal{X}^{\prime}}(k_{y}^{\prime}+1)+\sum_{y\in\mathcal{X}\backslash\mathcal{X}^{\prime}}k_{y}^{\prime}=\sum_{y\in\mathcal{X}}k_{y}^{\prime}+\sum_{y\in\mathcal{X}^{\prime}}1=\sum_{y\in\mathcal{X}}k_{y}^{\prime}+n-\sum_{y\in\mathcal{X}}k_{y}^{\prime}=n$
    and $k_{y}\in\{0,1,\dots,n\}$. We continue to define a probability measure $\mathbb{P}^{(n)}\in G^{(n)}$ through
    $\mathbb{P}^{(n)}(y):=\frac{k_{y}}{n}$
    Note that $\mathbb{P}^{(n)}(y)\in\left\{\frac{\lfloor n\cdot\mathbb{P}(y)\rfloor}{n},\frac{\lceil n\cdot\mathbb{P}(y)\rceil}{n}\right\}$, and hence
    (7.37) $\left|\mathbb{P}^{(n)}(y)-\mathbb{P}(y)\right|\leq\frac{1}{n}.$
    Without loss of generality assume $\mathbb{P}^{(n)}\neq\mathbb{P}$. Then, define the joint distribution $\pi\in{\Pi}(\mathbb{P}^{(n)},\mathbb{P})$ via
    $\pi(y,y^{\prime}):=\begin{cases}\min\{\mathbb{P}(y),\mathbb{P}^{(n)}(y)\}&\text{ if }y=y^{\prime},\\ \frac{1}{T}\max\{\mathbb{P}(y)-\mathbb{P}^{(n)}(y),0\}\cdot\max\{\mathbb{P}^{(n)}(y^{\prime})-\mathbb{P}(y^{\prime}),0\}&\text{ if }y\neq y^{\prime}.\end{cases}$
    for $T:=\sum_{y\in\mathcal{X}}\max\{\mathbb{P}^{(n)}(y)-\mathbb{P}(y),0\}$. With this definition and by using (7.37), we have for all $p\in\mathbb{N}$
    $W_{p}(\mathbb{P},\mathbb{P}^{(n)})$ $=\left(\min_{\widetilde{\pi}\in{\Pi}(\mathbb{P},\mathbb{P}^{(n)})}\sum_{x,y\in\mathcal{X}}\|y-y^{\prime}\|^{p}\,\widetilde{\pi}(y,y^{\prime})\right)^{1/p}$
    $\leq\left(\sum_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|^{p}\pi(y,y^{\prime})\right)^{1/p}$
    $=\left(\sum_{y,y^{\prime}\in\mathcal{X},y\neq y^{\prime}}\|y-y^{\prime}\|^{p}\pi(y,y^{\prime})\right)^{1/p}$
    $\leq\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot\left(\sum_{y,y^{\prime}\in\mathcal{X},y\neq y^{\prime}}\pi(y,y^{\prime})\right)^{1/p}$
    $\leq\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot\left(\sum_{y\in\mathcal{X}}|\mathbb{P}^{(n)}(y)-\mathbb{P}(y)|\right)^{1/p}\leq\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/p}\cdot\frac{1}{n^{1/p}}$
    where $|\mathcal{X}|$ denotes the number of states.
  • (ii)
    Let $n\in\mathbb{N}$. We define, given some $q:\mathcal{X}\times A\rightarrow\mathbb{R}$, the operator
    (7.38) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}^{(n)}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right],$
    and observe that by Lemma 7.1 (which is applicable as $\mathcal{P}^{(n)}$ is a finite ambiguity set, and by Lemma 7.2
    $\|Q^{(n)^{*}}-Q^{*}\|_{\infty}$ $=\|\mathcal{H}^{(n)}Q^{(n)^{*}}-\mathcal{H}Q^{*}\|_{\infty}$
    $\leq\|\mathcal{H}^{(n)}Q^{(n)^{*}}-\mathcal{H}^{(n)}Q^{*}+\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
    $\leq\alpha\|Q^{(n)^{*}}-Q^{*}\|_{\infty}+\|\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
    and thus
    (7.39) $\|Q^{(n)^{*}}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}.$
    Moreover for all $(x,a)\in\mathcal{X}\times A$ we have
    (7.40) $\mathcal{H}^{(n)}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)$
    $\min_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]-\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]$
    We first select a minimizing $\underline{\mathbb{P}}\in\mathcal{P}(x,a)$, i.e.,
    $\min_{\mathbb{P}\in\mathcal{P}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]=\mathbb{E}_{\underline{\mathbb{P}}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}Q^{*}(X_{1},b)\right]$
    whose existence is ensured by Assumption 4.1 (see [38, Theorem 2.7 (i) & Proposition 3.1]), and then, we obtain by Lemma 4.4 (i) some $\underline{\mathbb{P}}^{(n)}\in G^{(n)}$ with $W_{q}(\underline{\mathbb{P}}^{(n)},\underline{\mathbb{P}})\leq\frac{\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q}}{n^{1/q}}$. Then in case that $W_{q}(\underline{\mathbb{P}},\widehat{\mathbb{P}}(x,a))<\varepsilon$ we have
    $W_{q}(\underline{\mathbb{P}}^{(n)},\widehat{\mathbb{P}}(x,a))\leq W_{q}(\underline{\mathbb{P}}^{(n)},\underline{\mathbb{P}})+W_{q}(\underline{\mathbb{P}},\widehat{\mathbb{P}}(x,a))<\varepsilon$
    for $n$ large enough, and thus $\underline{\mathbb{P}}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ for $n$ large enough.
    If $\underline{\mathbb{P}}$ lies on the boundary of the Wasserstein-ball first consider the convex combination
    $\mathbb{Q}^{(n)}:=(1-1/n^{1/q})\underline{\mathbb{P}}+1/n^{1/q}\widehat{\mathbb{P}}(x,a)$
    which implies
    $W_{q}(\mathbb{Q}^{(n)},\widehat{\mathbb{P}}(x,a))\leq(1-\frac{1}{n^{1/q}})W_{q}(\mathbb{P},\widehat{\mathbb{P}}(x,a))=(1-\frac{1}{n^{1/q}})\varepsilon<\varepsilon$
    Hence $\mathbb{Q}^{(n)}$ is in the Wasserstein-ball, and as in the first case we obtain obtain by Lemma 4.4 (i) some $\underline{\mathbb{P}}^{(n)}\in G^{(n)}$ with $W_{q}(\underline{\mathbb{P}}^{(n)},\mathbb{Q}^{(n)})\leq\frac{\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q}}{n^{1/q}}$ lying in the Wasserstein-ball for $n$ large enough, implying
    (7.41) $W_{q}(\underline{\mathbb{P}},\underline{\mathbb{P}}^{(n)})\leq W_{q}(\underline{\mathbb{P}},\mathbb{Q}^{(n)})+W_{q}(\mathbb{Q}^{(n)},\underline{\mathbb{P}}^{(n)})\leq\frac{1}{n^{1/q}}\varepsilon+\frac{\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q}}{n^{1/q}}.$
    which holds true in both cases. Together with (7.40), we obtain
    $\mathcal{H}^{(n)}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)$
    $\leq(L_{r}+\alpha L_{V})\left\{\mathbb{E}_{\underline{\mathbb{P}}^{(n)}}\left[\frac{r(x,a,X_{1})+\alpha V^{*}(X_{1})}{(L_{r}+\alpha L_{V})}\right]-\mathbb{E}_{\underline{\mathbb{P}}}\left[\frac{r(x,a,X_{1})+\alpha V^{*}(X_{1})}{(L_{r}+\alpha L_{V})}\right]\right\}$
    $\leq(L_{r}+\alpha L_{V})\sup_{f:\mathcal{X}\rightarrow\mathbb{R},1-\text{Lipschitz}}\left\{\mathbb{E}_{\underline{\mathbb{P}}^{(n)}}\left[f(X_{1})\right]-\mathbb{E}_{\underline{\mathbb{P}}}\left[f(X_{1})\right]\right\}=(L_{r}+\alpha L_{V})W_{1}(\underline{\mathbb{P}},\underline{\mathbb{P}}^{(n)})$
    from the Kantorovich–Rubinstein duality. Moreover, $\mathcal{H}Q^{*}(x,a)-\mathcal{H}^{(n)}Q^{*}(x,a)\leq 0$ implying
    (7.42) $\left|\mathcal{H}^{(n)}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)\right|\leq(L_{r}+\alpha L_{V})W_{1}(\underline{\mathbb{P}},\underline{\mathbb{P}}^{(n)}).$
    By using $W_{1}\leq W_{q}$ (see [58, Remark 6]), and by (7.39), (7.41), and (7.42) we then obtain
    $\|Q^{(n)^{*}}-Q^{*}\|_{\infty}\leq\frac{(L_{r}+\alpha L_{V})W_{1}(\underline{\mathbb{P}},\underline{\mathbb{P}}^{(n)})}{1-\alpha}\leq\frac{(L_{r}+\alpha L_{V})(\varepsilon+\max_{y,y^{\prime}\in\mathcal{X}}\|y-y^{\prime}\|\cdot|\mathcal{X}|^{1/q})}{(1-\alpha)n^{1/q}}.$
Proof of Lemma 4.5.
Let $(x,a)\in\mathcal{X}\times A$. Let $n\in\mathbb{N}$ and consider the open cover of $\Theta(x,a)$.
$\left\{B_{1/n}(\theta)~|~\theta\in\Theta(x,a)\right\}\supseteq\Theta(x,a),$
where $B_{1/n}(\theta)$ denotes (open) balls centered at $\theta$ with radius $1/n$. By the compactness of $\Theta(x,a)$, there exists a finite subcover, i.e., there exist $\theta_{1}^{(n)},\dots,\theta_{m_{n}}^{(n)}\in\Theta(x,a)$ such that $\Theta(x,a)\subset\cup_{k=1}^{m_{n}}B_{1/n}(\theta_{k})$. Then, we define
$\Theta^{(n)}(x,a):=\{\theta_{1}^{(n)},\dots,\theta_{m_{n}}^{(n)}\}\subseteq\Theta(x,a)$
and $\mathcal{P}^{(n)}(x,a):=\mathcal{P}^{\Theta^{(n)}}(x,a)$. To show that Lemma 4.2 applies, we first note that Assumption 4.1 is implied by the assumed continuity of $\theta\mapsto\widehat{\mathbb{P}}(x,a,\theta)$, see [38, Proposition 3.2]. We pick some $\mathbb{P}\in\mathcal{P}(x,a)$ which, by definition of $\mathcal{P}$, can be represented as $\mathbb{P}=\widehat{\mathbb{P}}(x,a,\theta)$ for some $\theta\in\Theta(x,a)$. By construction of $\Theta^{(n)}$, there exists some $\theta^{(n)}\in\Theta^{(n)}(x,a)$ with $d(\theta,\theta^{(n)})\leq 1/n$, and the assertion of Lemma 4.2 follows by defining $\mathbb{P}^{(n)}=\widehat{\mathbb{P}}(x,a,\theta^{(n)})$ and since $\widehat{\mathbb{P}}(x,a,\theta^{(n)})\rightarrow\widehat{\mathbb{P}}(x,a,\theta)=\mathbb{P}$ in the Wasserstein-1 topology as $n\rightarrow\infty$ with the assumed continuity of $\theta\mapsto\widehat{\mathbb{P}}(x,a,\theta)$. ∎
Proof of Example 4.6.
We first show that [38, Assumption 2.2] is fulfilled for the weak topology (i.e., $p=0$ in the notation of [38]). This means we want to show that
$\mathcal{X}\times A\ni(x,a)\twoheadrightarrow\mathcal{P}(x,a):=\{\operatorname{Bin}(N,p),p\in[\underline{p}(x),\overline{p}(x)]\}$
is non-empty, compact-valued, and continuous.
The non-emptiness follows from requiring $\underline{p}(\cdot)\leq\overline{p}(\cdot)$.
For the compactness let $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}\subseteq\mathcal{P}(x,a)$, then we have for all $n\in\mathbb{N}$ the representation $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for some $p^{(n)}\in[\underline{p}(x),\overline{p}(x)]$. Hence, by the Bolzano–Weierstrass theorem there exists a convergent subsequence $p^{(n_{k})}\rightarrow p\in[\underline{p}(x),\overline{p}(x)]$ as $k\rightarrow\infty$ which also implies $\mathbb{P}^{(n_{k})}\rightarrow\mathbb{P}=\operatorname{Bin}(N,p)$ weakly.
For the continuity we use the characterization of upper and lower hemicontinuity provided in [2, Theorem 17.20 and Theorem 17.21]. For the upper hemicontinuity take a sequence $(x^{(n)},a^{(n)})\subseteq\mathcal{X}\times A$ with $(x^{(n)},a^{(n)})\rightarrow(x,a)\in\mathcal{X}\times A$ as $n\rightarrow\infty$, and consider a sequence $(\mathbb{P}^{(n)})_{n\in\mathbb{N}}$ with $\mathbb{P}^{(n)}\in\mathcal{P}(x^{(n)},a^{(n)})$ for all $n\in\mathbb{N}$. Then, we can represent $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for some $p^{(n)}\in[\underline{p}(x^{(n)}),\overline{p}(x^{(n)})]$ and for $n$ large enough $x^{(n)}=x$, hence $p^{(n)}\in[\underline{p}(x),\overline{p}(x)]$, i.e., $\mathbb{P}^{(n)}\in\mathcal{P}(x,a)$ showing the upper hemicontinuity. To show the lower hemicontinuity, take a sequence $(x^{(n)},a^{(n)})\subseteq\mathcal{X}\times A$ with $(x^{(n)},a^{(n)})\rightarrow(x,a)\in\mathcal{X}\times A$ and let $\mathbb{P}\in\mathcal{P}(x,a)$, then $\mathbb{P}=\operatorname{Bin}(N,p)$ for some $p=\lambda\underline{p}(x)+(1-\lambda)\overline{p}(x)$ with $\lambda\in[0,1]$. Define $\mathbb{P}^{(n)}=\operatorname{Bin}(N,p^{(n)})$ for $p^{(n)}=\lambda\underline{p}(x^{(n)})+(1-\lambda)\overline{p}(x^{(n)})$ implying $\mathbb{P}^{(n)}\in\mathcal{P}(x^{(n)},a^{(n)})$. Then, by definition $p^{(n)}\rightarrow p$ and hence $\operatorname{Bin}(N,p^{(n)})\rightarrow\operatorname{Bin}(N,p)$ as $n\rightarrow\infty$ weakly, showing the lower hemicontinuity.
Next, we show that $\mathcal{P}^{(n)}(x,a)$ fulfils the assumptions of Lemma 4.2. To this end, let $(x,a)\in\mathcal{X}\times A$ and $\mathbb{P}\in\mathcal{P}(x,a)$, i.e., $\mathbb{P}=\operatorname{Bin}(N,p)$ for some $p=\lambda\underline{p}(x)+(1-\lambda)\overline{p}(x)$ with $\lambda\in[0,1]$. Define $\mathbb{P}^{(n)}\in\mathcal{P}^{(n)}(x,a)$ by $\mathbb{P}^{(n)}:=\operatorname{Bin}(N,p_{i})$ where $i\in\{1,\dots,n\}$ such that $|\frac{i}{n}-\lambda|$ is minimal, implying that $p_{i}\rightarrow p$ as $n\rightarrow\infty$. Then, we have for all continuous and bounded functions $f:\mathcal{X}\rightarrow\mathbb{R}$ that
$\int_{\mathcal{X}}f(x){\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)$ $=\sum_{k=0}^{N}{N\choose k}p_{i}^{k}(1-p_{i})^{k}f(k)\rightarrow\sum_{k=0}^{N}{N\choose k}p^{k}(1-p)^{k}f(k)=\int_{\mathcal{X}}f(x){\mathop{}\!\mathrm{d}}\mathbb{P}(x)$
as $n\rightarrow\infty$, i.e., $\mathbb{P}^{(n)}\rightarrow\mathbb{P}$ weakly. Analogously, we obtain $\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}^{(n)}(x)\rightarrow\int_{\mathcal{X}}|x|{\mathop{}\!\mathrm{d}}\mathbb{P}(x)$. ∎
Proof of Corollary 4.7.
We define
(7.43) $\mathcal{X}\times A\ni(x,a)\mapsto\mathcal{H}^{(n)}q(x,a):=\min_{\mathbb{P}\in\mathcal{P}^{(n)}(x,a)}\mathbb{E}_{\mathbb{P}}\left[r(x,a,X_{1})+\alpha\max_{b\in A}q(X_{1},b)\right],$
and observe that by Lemma 7.1 and Lemma 7.2
$\|Q^{(n)^{*}}-Q^{*}\|_{\infty}$ $=\|\mathcal{H}^{(n)}Q^{(n)^{*}}-\mathcal{H}Q^{*}\|_{\infty}$
$\leq\|\mathcal{H}^{(n)}Q^{(n)^{*}}-\mathcal{H}^{(n)}Q^{*}+\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
$\leq\alpha\|Q^{(n)^{*}}-Q^{*}\|_{\infty}+\|\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}$
and thus
(7.44) $\|Q^{(n)^{*}}-Q^{*}\|_{\infty}\leq\frac{\|\mathcal{H}^{(n)}Q^{*}-\mathcal{H}Q^{*}\|_{\infty}}{1-\alpha}.$
Let $(x,a)\in\mathcal{X}\times A$. For any $\theta\in\Theta(x,a)$ choose $\theta^{(n)}\in\Theta^{(n)}(x,a)$ with $\|\theta-\theta^{(n)}\|\leq\varepsilon_{n}$ where $\varepsilon_{n}$ is defined in (4.8). By (4.7) we get $W_{1}(\widehat{\mathbb{P}}(x,a,\theta),\widehat{\mathbb{P}}(x,a,\theta^{(n)}))\leq L_{\theta}\varepsilon_{n}$. Using the Kantorovich–Rubinstein duality and the Lipschitz bounds for $y\mapsto r(x,a,y)$ and $y\mapsto V^{\ast}(y)$ yields
$\mathcal{H}^{(n)}Q^{*}(x,a)-\mathcal{H}Q^{*}(x,a)$
$\leq(L_{r}+\alpha L_{V})\left\{\mathbb{E}_{\widehat{\mathbb{P}}(x,a,\theta^{(n)})}\left[\frac{r(x,a,X_{1})+\alpha V(X_{1})}{(L_{r}+\alpha L_{V})}\right]-\mathbb{E}_{\widehat{\mathbb{P}}(x,a,\theta)}\left[\frac{r(x,a,X_{1})+\alpha V(X_{1})}{(L_{r}+\alpha L_{V})}\right]\right\}$
$\leq(L_{r}+\alpha L_{V})\sup_{f:\mathcal{X}\rightarrow\mathbb{R},1-\text{Lipschitz}}\left\{\mathbb{E}_{\widehat{\mathbb{P}}(x,a,\theta^{(n)})}\left[f(X_{1})\right]-\mathbb{E}_{\widehat{\mathbb{P}}(x,a,\theta)}\left[f(X_{1})\right]\right\}$
$=(L_{r}+\alpha L_{V})W_{1}\left(\widehat{\mathbb{P}}(x,a,\theta^{(n)}),\widehat{\mathbb{P}}(x,a,\theta)\right)$
$\leq(L_{r}+\alpha L_{V})L_{\theta}\varepsilon_{n}$
where we have chosen $\theta$ as the minimizing parameter of $\mathcal{H}Q^{*}(x,a)$ whose existence is ensured by the assumed continuity of $\theta\mapsto\widehat{\mathbb{P}}(x,a,\theta)$ (see [38, Theorem 2.7 (i) & Proposition 3.2]). Together with (7.44) we get the assertion. ∎
Acknowledgements
J.S. gratefully acknowledges financial support by the NUS Start-Up Grant Tackling model uncertainty in Finance with machine learning.